### v2

# Heavy neutral bosons and dark matter in the 3-3-1 model with axionlike particle

Abstract
We consider heavy neutral bosons in the 3-3-1 model with axionlike particles (331ALP), including the Higgs boson and the $Z^{\prime}$ boson which are outside the standard model (SM). Based on gluon-gluon fusion at the LHC, we investigate the signals of cross-sections in the parameter space region satisfying the current experimental limits of lepton flavor violating decay, including processes involving both charged leptons and Higgs boson, and provide predictions of $m_{h_{2}}\geq 600\penalty 10000\ \mathrm{GeV}$. A new gauge boson, labeled as $Z^{\prime}$, is predicted $m_{Z^{\prime}}\geq 5.1\penalty 10000\ \mathrm{TeV}$ based on the search for high-mass dilepton resonances at ATLAS and CMS. We consider the stability of odd-$Z_{2}$ particles, with $Z_{2}$ is assumed a residual symmetry after spontaneous symmetry breaking stages, to point out dark matter candidates in the model. Investigating the relic density of dark matter within experimentally permissible limits, we established a relationship between the mass of dark matter and the breaking scale of axion.
dark matter, heavy neutral bosons
I Introduction
The hypothesis of lepton flavor violating (LFV) processes is increasingly strengthened through increasingly accurate experimental data. The neutral lepton part is trusted with confirmation of the mass and oscillations of neutrinos Patrignani:2016xqp ; Tanabashi:2018oca , and the charged lepton part is also supported with given limits on branching ratios as ATLAS:2019erb ; Zyla:2020zbs ; BaBar:2009hkt ; MEG:2016leq :
$\mathrm{Br}(\mu\rightarrow e\gamma)\leq 4.2\times 10^{-13},\,\mathrm{Br}(\tau\rightarrow e\gamma)\leq 3.3\times 10^{-8},\,\mathrm{Br}(\tau\rightarrow\mu\gamma)\leq 4.4\times 10^{-8},$ (1)
and the lepton flavor violating decays of SM-like Higgs boson (LFVHD) also have experimental limits given as CMS:2018ipm ; ATLAS:2019pmk ; ATLAS:2019old .
$\mathrm{Br}(h\rightarrow\mu\tau)\leq 10^{-3},\,\mathrm{Br}(h\rightarrow\tau e)\leq 10^{-3},\,\mathrm{Br}(h\rightarrow\mu e)\leq 6.1\times 10^{-5}.$ (2)
The LFV processes are of greater interest in the models beyond the standard model (BSM), where are abundant sources of lepton flavor violation. They may appear alongside mass generation mechanisms for neutrinos, including two main types: effective operators Mizukoshi:2010ky ; Dias:2005yh ; Hung:2022vqx ; Hung:2022kmv ; PhysRevD.47.2918 and seesaw mechanisms Boucenna:2015zwa ; Hernandez:2013hea ; Nguyen:2018rlb ; Hung:2021fzb ; Ferreira:2019qpf ; CarcamoHernandez:2019pmy ; Catano:2012kw ; Hernandez:2014lpa ; Dias:2012xp . Both these types yielded good results regarding LFVHD, even showing that $Br(h\rightarrow\mu\tau)$ can reach $\mathcal{O}(10^{-4})$ Hue:2015fbb ; Thuc:2016qva ; Zhang:2015csm ; Herrero-Garcia:2017xdu . Theoretical results have shown that the signals of LFVHD are much smaller than the experimental limit at the LHC Blankenburg:2012ex , even with the contribution of Majorana neutrinos Pilaftsis:1992st . One publication also pointed out that the LFVHD signal at one loop order is always less than $10^{-4}$, because it is suppressed by a loop factor and constraints from the lepton flavor violating of charged lepton (cLFV) Herrero-Garcia:2016uab . Furthermore, some publications have indicated a parameter space region that satisfies the experimental limits of both cLFV and LFVHD, which is well-suited for investigating other phenomena, for example, explaining the supplemental part for the W boson mass through LFV processes Duy:2024txy , predicting the signal $(g-2)_{\mu}$ in the parameter space region satisfying the experimental limits of LFVHD Hong:2024yhk , studying $Z\rightarrow e_{a}e_{b}$ associated with other LFV processes Hong:2024swk .
Recently, the experimental LHC based on gluon-gluon fusion at a center of mass energy $\sqrt{s}=13\penalty 10000\ \mathrm{TeV}$ have yielded very noteworthy results. The search for a heavy neutral Higgs boson in the range of $200\to 900\penalty 10000\ \mathrm{GeV}$, based on the $\mu\tau$ decay mode, performed by CMS, using a CMS detector with an intergrated luminosity of $35.9\penalty 10000\ \mathrm{fb}^{-1}$, has shown upper limits on production cross section multiplied by the branching ratio to be in the range of $51.9(57.4)\penalty 10000\ \mathrm{fb}$ to $1.6(2.1)\penalty 10000\ \mathrm{fb}$ at $95\%\penalty 10000\ \mathrm{CL}$ CMS:2019pex . Furthermore, the search for high-mass resonances in the range of $250\to 6000\penalty 10000\ \mathrm{GeV}$ has also been carried out by ATLAS ($139\penalty 10000\ \mathrm{fb}^{-1}$) and CMS($140\penalty 10000\ \mathrm{fb}^{-1}$). The upper limits of the cross-section times branching ratio are given at $95\%\penalty 10000\ \mathrm{CL}$, which makes predicting the masses of heavy neutral bosons even more exciting ATLAS:2019erb ; CMS:2021ctt .
The 331ALP possesses many interesting properties; by adding the scalar gauge singlet, named as $\phi$, the axion generation mechanism is naturally resolved. The masses of other particles are generated in a complete way thanks to the symmetry group $Z(11)\otimes Z(2)$. The oscillations of ordinary neutrinos are explained by $N_{aR}$ alp331 . However, LFV processes, the masses of heavy neutral bosons, and dark matter have yet to be considered. In this work, we will address those issues.
The paper is organized as follows. In the next section, we review the model and give masses spectrum of gauge and Higgs bosons. We calculate the Feynman rules and analytic formulas for cLFV and LFVHDs in Section III. $Z^{\prime}$ boson and dark mater are discussed in Section IV and Section V. Conclusions are in Section VI. Finally, we provide Appendix A,B to calculate the amplitude and show numerical results of $h_{1,2}\rightarrow\mu\tau$ decays.
II Review of the model
II.1 Particle content and discrete symmetries
The fermion sector of the model and their $\mbox{SU(3)}_{C}\times\mbox{SU(3)}_{L}\times\mbox{U(1)}_{X}$ assignments are:
$\psi_{aL}=\left(\nu_{a},l_{a},(\nu_{aR})^{c}\right)_{L}^{T}\sim\left(1,3,-1/3\right),\hskip 8.5359ptl_{aR}\sim\left(1,1,-1\right),\penalty 10000\ N_{aR}\sim(1,1,0),$
$Q_{3L}=\left(u_{3},d_{3},U_{3}\right)_{L}^{T}\sim\left(3,3,1/3\right),\penalty 10000\ Q_{nL}=\left(d_{n},-u_{n},D_{n}\right)_{L}^{T}\sim\left(3,3^{\ast},0\right),$
$u_{aR}\sim\left(3,1,2/3\right),\penalty 10000\ U_{3R}\sim\left(3,1,2/3\right),\hskip 8.5359ptd_{aR}\sim(3,1,-1/3),\penalty 10000\ D_{nR}\sim(3,1,-1/3),$ (3)
where $n=1,2$ and $a=\{n,3\}$ are family indices. The $U_{3}$ and $D_{n}$ are exotic quarks with ordinary electric charges, whereas $N_{aR}$ and $\nu_{aR}$ are right-handed Majorana and Dirac neutrinos, respectively.
The third component of the $SU(3)_{L}$ triplet leptons are $(\nu_{aR})^{c}$, so the simplest way to give them mass is to use the seesaw mechanism through a combination with $N_{aR}$. To perform this, the $Z_{11}$ group with charged of the form $\omega_{m}\equiv e^{\frac{i2\pi m}{11}},m=0,\pm 1\cdots\pm 5$ is introduced. Furthermore, 3-3-1 models with $\beta=\pm\frac{1}{\sqrt{3}}$ often appear the $SU(3)_{L}$ triplet scalars which have the same quantum numbers, it leads to difficulties in deriving physical states due to self-coupling terms of these fields. To avoid this, the $Z_{2}$ symmetry is introduced into the model, under the transformation of this group, the odd fields are given as alp331 ,
$(\eta\,,\rho\,,u_{aR}\,,d_{nR},l_{aR},N_{aR})\,\,\,\,\rightarrow\,\,\,\,-\,(\eta\,,\rho\,,u_{aR}\,,d_{nR},l_{aR},N_{aR})\,.$ (4)
We need three $SU(3)_{L}$ triplet scalars, $\eta,\rho,\chi$, to generate mass for gauge bosons and fermions as usual in 3-3-1 models with right-handed neutrinos. However, in this case, we require an additional neutral $SU(3)_{L}$ singlet scalar, $\phi$, to generate mass for exotic neutrinos. Therefore, the multiplet scalars and their VeVs are given as follows alp331 ; Long:2024rdq :
$\eta=\left(\begin{array}[]{c}\eta_{1}^{0}\\ \eta_{2}^{-}\\ \eta_{3}^{0}\end{array}\right)\sim\left(1,3,-\frac{1}{3}\right),\,\chi=\left(\begin{array}[]{c}\chi_{1}^{0}\\ \chi_{2}^{-}\\ \chi_{3}^{0}\end{array}\right)\sim\left(1,3,-\frac{1}{3}\right),\,\rho=\left(\begin{array}[]{c}\rho_{1}^{+}\\ \rho_{2}^{0}\\ \rho_{3}^{+}\end{array}\right)\sim\left(1,3,\frac{2}{3}\right),\,\phi\sim(1,1,0)\,$, (14)
and
$\langle\eta\rangle=\left(\begin{array}[]{c}\frac{v_{\eta}}{\sqrt{2}}\\ 0\\ 0\end{array}\right),\,\langle\chi\rangle=\left(\begin{array}[]{c}0\\ 0\\ \frac{v_{\chi}}{\sqrt{2}}\end{array}\right),\,\langle\rho\rangle=\left(\begin{array}[]{c}0\\ \frac{v_{\rho}}{\sqrt{2}}\\ 0\end{array}\right),\,\langle\phi\rangle=\frac{v_{\phi}}{\sqrt{2}}\,.$ (24)
By introducing the discrete groups based on the above discussions, the full gauge symmetry group of the model is $\mbox{SU(3)}_{C}\otimes\mbox{SU(3)}_{L}\otimes\mbox{U(1)}_{X}\otimes Z_{11}\otimes Z_{2}$ and under the transformation of this group the particle assignments are summarized in table 1.
$Q_{nL}$ $Q_{3L}$ $u_{aR}$ $d_{aR}$ $U_{3R}$ $D_{nR}$ $\psi_{aL}$ $l_{aR}$ $N_{aR}$ $\eta$ $\chi$ $\rho$ $\phi$
$SU(3)_{C}$ $\mathbf{3}$ $\mathbf{3}$ $\mathbf{3}$ $\mathbf{3}$ $\mathbf{3}$ $\mathbf{3}$ $\mathbf{1}$ $\mathbf{1}$ $\mathbf{1}$ $\mathbf{1}$ $\mathbf{1}$ $\mathbf{1}$ $\mathbf{1}$
$SU(3)_{L}$ $\overline{\mathbf{3}}$ $\mathbf{3}$ $\mathbf{1}$ $\mathbf{1}$ $\mathbf{1}$ $\mathbf{1}$ $\mathbf{3}$ $\mathbf{1}$ $\mathbf{1}$ $\mathbf{3}$ $\mathbf{3}$ $\mathbf{3}$ $\mathbf{1}$
$U(1)_{X}$ 0 $\frac{1}{3}$ $\frac{2}{3}$ $\frac{1}{3}$ $\frac{2}{3}$ $\frac{1}{3}$ $\frac{1}{3}$ 1-1 0 $\frac{1}{3}$ $\frac{1}{3}$ $\frac{2}{3}$ 0
$Z_{11}$ $\omega^{-1}_{4}$ $\omega_{0}$ $\omega_{5}$ $\omega_{2}$ $\omega_{3}$ $\omega_{4}$ $\omega_{1}$ $\omega_{3}$ $\omega^{-1}_{5}$ $\omega^{-1}_{5}$ $\omega^{-1}_{3}$ $\omega^{-1}_{2}$ $\omega^{-1}_{1}$
$Z_{2}$ 11 11 1-1 1-1 11 11 11 1-1 1-1 1-1 11 1-1 11
Table 1: The $SU(3)_{C}\otimes SU(3)_{L}\otimes U(1)_{X}\otimes Z_{11}\otimes Z_{2}$ charge assignments of spectrum of the particles in the model with $a=1,2,3$ and $n=1,2$.
From table 1, one can derive the Yukawa interactions of the model with two parts corresponding to quarks and leptons, in which the terms that do not preserve the charge of the symmetry gauge group are eliminated. The Yukawa couplings invariant under the $SU(3)_{C}\otimes SU(3)_{L}\otimes U(1)_{X}\otimes Z_{11}\otimes Z_{2}$ symmetry, arise alp331 :
$-\mathcal{L}^{Y}_{f}$ $=$ $-\mathcal{L}^{Y}_{q}-\mathcal{L}^{Y}_{\ell},$
$-\mathcal{L}^{Y}_{q}$ $=$ $y_{1}\bar{Q}_{3L}U_{R}\chi+\sum\limits_{n,p=1}^{2}\left(y_{2}\right)_{np}\bar{Q}_{nL}D_{pR}\chi^{\ast}$
$+\sum\limits_{i=1}^{3}\left(y_{3}\right)_{3a}\bar{Q}_{3L}u_{aR}\eta+\sum\limits_{n=1}^{2}\sum\limits_{i=1}^{3}\left(y_{4}\right)_{ni}\bar{Q}_{nL}d_{iR}\eta^{\ast}$
$+\sum\limits_{i=1}^{3}\left(y_{5}\right)_{3i}\bar{Q}_{3L}d_{iR}\rho+\sum\limits_{n=1}^{2}\sum\limits_{i=1}^{3}\left(y_{6}\right)_{ni}\bar{Q}_{nL}u_{iR}\rho^{\ast},$
$-\mathcal{L}^{Y}_{\ell}$ $=$ $\sum\limits_{i=1}^{3}\sum\limits_{j=1}^{3}g_{ij}\bar{\psi}_{iL}l_{jR}\rho+\sum\limits_{i=1}^{3}\sum\limits_{j=1}^{3}\left(y_{\nu}^{D}\right)_{ij}\bar{\psi}_{iL}\eta N_{jR}$ (25)
$+\sum\limits_{i=1}^{3}\sum\limits_{j=1}^{3}\left(y_{N}\right)_{ij}\phi\bar{N}_{iR}^{C}N_{jR}+\mbox{H.c.}$
We can easily see that the Lagrangian in Eq.(25) is conserved charge for all elements of the gauge group, including the $Z_{2}$ group as usual. It is emphasized that the transformation under the $Z_{2}$ as in Eq. (4) is different from the one given in Ref. jpf where $\chi$ is odd. So, $\chi$ only couple with exotic quarks and is responsible for giving mass to particles outside the standard model except $N_{aR}$. The masses of $N_{aR}$ are generated through $v_{\phi}$ of the gauge singlet $\phi$. Furthermore, $v_{\phi}$ also breaks the $Z_{11}$ symmetry to helps the axion-like particle and an inflation arise. Two remaining $SU(3)_{L}$ triplet scalars $\eta$ and $\rho$ generate masses for SM particles as in the usual 3-3-1 models. To fully illustrate this, we present the following scheme of symmetry breaking:
$SU(3)_{C}\otimes SU(3)_{L}$ $\otimes$ $U(1)_{X}\otimes Z(11)\otimes Z_{2}$
$\downarrow$ $v_{\phi}$
$SU(3)_{C}\otimes SU(3)_{L}$ $\otimes$ $U(1)_{X}\otimes Z_{2}$
$\downarrow$ $v_{\chi}$
$SU(3)_{C}\otimes SU(2)_{L}$ $\otimes$ $U(1)_{Y}\otimes Z_{2}$
$\downarrow$ $v_{\eta},\,v_{\rho}$
$SU(3)_{C}$ $\otimes$ $U(1)_{Q}\otimes Z_{2}$ (26)
Clearly, $Z_{2}$ is a residual symmetry after the spontaneous symmetry-breaking stages, and it plays a crucial role in identifying stable particles in the model, which may include dark matter candidates.
II.2 Gauge bosons
We consider the structure of gauge bosons from the covariant derivative with nine electroweak gauge bosons of the $SU(3)_{L}\times U(1)_{X}$ symmetry. The kinetic terms of scalar fields are given by:
$\mathcal{L}_{S}=\sum_{S=\chi,\eta,\rho,\phi}(D^{\mu}S)^{\dagger}D_{\mu}S\,$, (27)
the covariant derivative is as follow,
$D_{\mu}\equiv\partial_{\mu}-igT^{a}W^{a}_{\mu}-ig_{X}XT^{9}B_{\mu}\equiv\partial_{\mu}-iP_{\mu},$ (28)
where $T^{9}=\frac{I_{3\times 3}}{\sqrt{6}}$ with $I_{3\times 3}$ is the $3\times 3$ identity matrix and $g$, $g_{X}$ are gauge couplings of the two groups $SU(3)_{L}$ and $U(1)_{X}$, respectively. Furthermore, the matrix $W^{a}T^{a}$, where $T^{a}=\frac{\lambda_{a}}{2}$ corresponds to a triplet representation, is written as follows:
$P_{\mu}=\frac{g}{2}\left(\begin{array}[]{ccc}W^{3}_{\mu}+\frac{1}{\sqrt{3}}W^{8}_{\mu}+t\sqrt{\frac{2}{3}}XB_{\mu}&\sqrt{2}W^{+}_{\mu}&\sqrt{2}X^{0}_{\mu}\\ \sqrt{2}W^{-}_{\mu}&-W^{3}_{\mu}+\frac{1}{\sqrt{3}}W^{8}_{\mu}+t\sqrt{\frac{2}{3}}XB_{\mu}&\sqrt{2}V^{-}_{\mu}\\ \sqrt{2}X^{0*}_{\mu}&\sqrt{2}YV^{+}_{\mu}&-\frac{2}{\sqrt{3}}W^{8}_{\mu}+t\sqrt{\frac{2}{3}}XB_{\mu}\\ \end{array}\right),$ (32)
in which we have denoted $t=\frac{g_{X}}{g}$ and defined the mass eigenstates of the charged gauge bosons as:
$W^{\pm}_{\mu}$ $=$ $\frac{1}{\sqrt{2}}\left(W^{1}_{\mu}\mp iW^{2}_{\mu}\right),\hskip 8.5359ptV^{\pm}_{\mu}=\frac{1}{\sqrt{2}}\left(W^{6}_{\mu}\pm iW^{7}_{\mu}\right),$
$X^{0}_{\mu}$ $=$ $\frac{1}{\sqrt{2}}\left(W^{4}_{\mu}-iW^{5}_{\mu}\right),\hskip 8.5359ptX^{0*}_{\mu}=\frac{1}{\sqrt{2}}\left(W^{4}_{\mu}+iW^{5}_{\mu}\right).$ (33)
After spontaneous symmetry breaking (SSB), the mass spectrum of the gauge bosons arises from the following kinertic terms,
$\mathcal{L}_{mass}=\sum_{S=\chi,\eta,\rho}(D^{\mu}\langle S\rangle)^{\dagger}(D_{\mu}\langle S\rangle)\,.$ (34)
The $W$ boson which is identical to that in the SM, while the charged and bileptonbilepton gauge bosons $(X,Y)$ get masses as below:
$m^{2}_{W}=\frac{g^{2}}{4}(v_{\eta}^{2}+v_{\rho}^{2}),\penalty 10000\ m^{2}_{X^{0}}=\frac{g^{2}}{4}(v_{\chi}^{2}+v_{\eta}^{2}),\hskip 8.5359ptm^{2}_{V}=\frac{g^{2}}{4}(v_{\chi}^{2}+v_{\rho}^{2})\,$ (35)
and their masses splitting with $W$ boson is il
$|m^{2}_{V}-m^{2}_{X^{0}}|<m^{2}_{W}\,.$
From Eq.(35), it follows
$v_{\eta}^{2}+v_{\rho}^{2}=v^{2}=246^{2}\,\textrm{GeV}^{2}\,.$ (36)
In the basis $(W^{3}_{\mu},W^{8}_{\mu},B_{\mu})$, the squared mass mixing matrix of neutral gauge bosons is
$\mathcal{M}^{2}_{NG}=\left(\begin{array}[]{ccc}\frac{g^{2}}{4}(v_{\eta}^{2}+v_{\rho}^{2})&\frac{g^{2}}{4\sqrt{3}}(v_{\eta}^{2}-v_{\rho}^{2})&-\frac{gg_{X}}{6\sqrt{6}}(v_{\eta}^{2}+2v_{\rho}^{2})\\ &\frac{g^{2}}{12}(v_{\eta}^{2}+v_{\rho}^{2}+4v_{\chi}^{2})&\frac{gg_{X}}{18\sqrt{2}}(2v_{\rho}^{2}+2v_{\chi}^{2}-v_{\eta}^{2})\\ &&\frac{g_{X}^{2}}{54}(v_{\eta}^{2}+4v_{\rho}^{2}+v_{\chi}^{2})\end{array}\right)\,.$ (40)
The matrix $\mathcal{M}^{2}_{NG}$ has three eigenvalues which belong to photon AA and $Z$, $Z^{\prime}$ bosons, respectively:
$m^{2}_{A}$ $=$ $0\,$,
$m^{2}_{Z}$ $=$ $\frac{g^{2}(9g^{2}+2g_{X}^{2})(v_{\eta}^{2}v_{\rho}^{2}+v_{\rho}^{2}v_{\chi}^{2}+v_{\chi}^{2}v_{\eta}^{2})}{2\left(18g^{2}(v_{\eta}^{2}+v_{\rho}^{2}+v_{\chi}^{2})+g_{X}^{2}(v_{\eta}^{2}+4v_{\rho}^{2}+v_{\chi}^{2})\right)}\,$,
$m^{2}_{Z^{\prime}}$ $=$ $\frac{g^{2}}{3}(v_{\eta}^{2}+v_{\rho}^{2}+v_{\chi}^{2})+\frac{g_{X}^{2}}{54}(v_{\eta}^{2}+4v_{\rho}^{2}+v_{\chi}^{2})-m^{2}_{Z}\,.$ (41)
The physical states of these neutral gauge boson are defined as below:
$\left(\begin{array}[]{c}A_{\mu}\\ Z_{\mu}\\ Z^{\prime}_{\mu}\end{array}\right)=\left(\begin{array}[]{ccc}c_{\beta_{2}}&c_{\beta_{1}}s_{\beta_{2}}&s_{\beta_{1}}s_{\beta_{2}}\\ -s_{\beta_{2}}&c_{\beta_{1}}c_{\beta_{2}}&c_{\beta_{2}}s_{\beta_{1}}\\ 0&-s_{\beta_{1}}&c_{\beta_{1}}\\ \end{array}\right)\left(\begin{array}[]{c}W^{3}_{\mu}\\ W^{8}_{\mu}\\ B_{\mu}\end{array}\right)\,$, (51)
where, we denote $c_{\varphi}=\cos\varphi$, $s_{\varphi}=\sin\varphi$, $t_{\varphi}=\tan\varphi$ and the mixing angles are given by:
$t_{2\beta_{1}}$ $=$ $\frac{6\sqrt{2}t_{W}(v_{\eta}^{2}-2v^{2}_{\rho}-2v^{2}_{\chi})}{2t_{W}^{2}(v^{2}_{\eta}+4v^{2}_{\rho}+v^{2}_{\chi})-9(v^{2}_{\eta}+v^{2}_{\rho}+4v^{2}_{\chi})}\,$,
$t_{2\beta_{2}}$ $=$ $-\frac{3\sqrt{6}c_{W}\left(B\left(v_{\eta}^{2}-v_{\rho}^{2}\right)+2\tan^{2}\theta_{W}\left(v_{\eta}^{4}-3v_{\eta}^{2}v_{\rho}^{2}-5v^{2}_{\eta}v_{\chi}^{2}-7v_{\rho}^{2}v_{\chi}^{2}-4v_{\rho}^{4}\right)\right)}{\left(v_{\eta}^{2}-2v_{\rho}^{2}-2v_{\chi}^{2}\right)\left(A-2t^{2}_{W}\left(v_{\eta}^{2}+4v_{\rho}^{2}+v_{\chi}^{2}\right)+45\left(v_{\eta}^{2}+v_{\rho}^{2}\right)-36v_{\chi}^{2}\right)\sqrt{1+\frac{\left(Bct_{W}-2t_{W}\left(v_{\eta}^{2}+4v_{\rho}^{2}+v_{\chi}^{2}\right)\right)^{2}}{72\left(v_{\eta}^{2}-2v_{\rho}^{2}-v_{\chi}^{2}\right)^{2}}}}\,$,
with $s_{w}\equiv\sin\theta_{W}$, $\theta_{W}$ is Weinberg angle as known, and
$A$ $=$ $\sqrt{\left(2t^{2}_{W}\left(v_{\eta}^{2}+4v_{\rho}^{2}+v_{\chi}^{2}\right)+9\left(v_{\eta}^{2}+v_{\rho}^{2}+4v_{\chi}^{2}\right)\right)^{2}-18(v_{\eta}^{2}v_{\rho}^{2}+v_{\eta}^{2}v_{\chi}^{2}+v_{\rho}^{2}v_{\chi}^{2})}\,$,
$B$ $=$ $9\left(v_{\eta}^{2}+v_{\rho}^{2}+4v_{\chi}^{2}\right)-A\,.$ (53)
In the limits $v_{\chi}\gg v_{\eta},v_{\rho}$, the mixing angle $\beta_{2}$ is very tiny, approximates zero, while the mixing angle $\beta_{1}$ is redefined as:
$t_{2\beta_{1}}\approx\frac{6\sqrt{2}t_{W}}{t^{2}_{W}-18}\,.$ (54)
II.3 The scalar sector
The scalar potential of 331ALP has the form alp331 :
$V$ $=$ $\mu^{2}_{\phi}\phi^{*}\phi+\mu_{\chi}^{2}\chi^{\dagger}\chi+\mu_{\rho}^{2}\rho^{\dagger}\rho+\mu_{\eta}^{2}\eta^{\dagger}\eta+\lambda_{1}(\chi^{\dagger}\chi)^{2}+\lambda_{2}(\eta^{\dagger}\eta)^{2}$ (55)
$+\lambda_{3}(\rho^{\dagger}\rho)^{2}+\lambda_{4}(\chi^{\dagger}\chi)(\eta^{\dagger}\eta)+\lambda_{5}(\chi^{\dagger}\chi)(\rho^{\dagger}\rho)+\lambda_{6}(\eta^{\dagger}\eta)(\rho^{\dagger}\rho)$
$+\lambda_{7}(\chi^{\dagger}\eta)(\eta^{\dagger}\chi)+\lambda_{8}(\chi^{\dagger}\rho)(\rho^{\dagger}\chi)+\lambda_{9}(\eta^{\dagger}\rho)(\rho^{\dagger}\eta)$
$+\lambda_{10}(\phi^{*}\phi)^{2}+\lambda_{11}(\phi^{*}\phi)(\chi^{\dagger}\chi)+\lambda_{12}(\phi^{*}\phi)(\rho^{\dagger}\rho)$
$+\lambda_{13}(\phi^{*}\phi)(\eta^{\dagger}\eta)+\left(\lambda_{\phi}\epsilon^{ijk}\eta_{i}\rho_{j}\chi_{k}\phi+H.c.\right).$
These scalar fields should be expanded around their VEVs,
$\rho_{2}^{0}$ $=$ $\frac{1}{\sqrt{2}}(v_{\rho}+R_{\rho}+iI_{\rho})\,,\hskip 8.5359pt\eta_{1}^{0}=\frac{1}{\sqrt{2}}(v_{\eta}+R^{1}_{\eta}+iI^{1}_{\eta})\,$,
$\chi_{3}^{0}$ $=$ $\frac{1}{\sqrt{2}}(v_{\chi}+R^{3}_{\chi}+iI^{3}_{\chi})\,,\hskip 8.5359pt\phi=\frac{1}{\sqrt{2}}(v_{\phi}+R_{\phi}+iI_{\phi})\,.$ (56)
Substituting (56) into (55) leads to the following minimum potential conditions,
$\mu^{2}_{\rho}+\lambda_{3}v_{\rho}^{2}+\frac{\lambda_{5}}{2}v_{\chi}^{2}+\frac{\lambda_{6}}{2}v_{\eta}^{2}+\frac{\lambda_{12}}{2}v_{\phi}^{2}+\frac{A}{2v_{\rho}^{2}}$ $=$ $0\,$,
$\mu^{2}_{\eta}+\lambda_{2}v_{\eta}^{2}+\frac{\lambda_{4}}{2}v_{\chi}^{2}+\frac{\lambda_{6}}{2}v_{\rho}^{2}+\frac{\lambda_{13}}{2}v_{\phi}^{2}+\frac{A}{2v_{\eta}^{2}}$ $=$ $0\,$,
$\mu^{2}_{\chi}+\lambda_{1}v_{\chi}^{2}+\frac{\lambda_{4}}{2}v_{\eta}^{2}+\frac{\lambda_{5}}{2}v_{\rho}^{2}+\frac{\lambda_{11}}{2}v_{\phi}^{2}+\frac{A}{2v_{\chi}^{2}}$ $=$ $0\,$,
$\mu^{2}_{\phi}+\lambda_{10}v_{\phi}^{2}+\frac{\lambda_{11}}{2}v_{\chi}^{2}+\frac{\lambda_{12}}{2}v_{\rho}^{2}+\frac{\lambda_{13}}{2}v_{\eta}^{2}+\frac{A}{2v_{\phi}^{2}}$ $=$ $0\,$, (57)
where $\mathcal{A}\equiv\lambda_{\phi}v_{\phi}v_{\chi}v_{\eta}v_{\rho}$.
The Higgs potential in Eq.(55) gives the mass mixing matrix of the charged scalar fields, diagonalizing it allows to give the physical state and masses of the charged Higgs bosons. The result obtained is similar to Ref.alp331 ,
$m^{2}_{H^{\pm}_{1}}=-\frac{v^{2}}{2}\left(\frac{\mathcal{A}}{v_{\rho}^{2}v_{\eta}^{2}}-\lambda_{9}\right)\,,\penalty 10000\ m^{2}_{H^{\pm}_{2}}=-\frac{(v_{\chi}^{2}+v_{\rho}^{2})}{2}\left(\frac{\mathcal{A}}{v_{\chi}^{2}v_{\rho}^{2}}-\lambda_{8}\right)\,$, (58)
with conditions $\lambda_{9}>\lambda_{\phi}\frac{v_{\phi}v_{\chi}}{v_{\rho}v_{\eta}}$, $\lambda_{8}>\lambda_{\phi}\frac{v_{\phi}v_{\eta}}{v_{\chi}v_{\rho}}$ and
$\left(\begin{array}[]{c}G_{W}^{\pm}\\ H^{\pm}_{1}\\ \end{array}\right)=\left(\begin{array}[]{cc}c_{\alpha}&-s_{\alpha}\\ s_{\alpha}&c_{\alpha}\\ \end{array}\right)\left(\begin{array}[]{c}\rho_{1}^{\pm}\\ \eta_{2}^{\pm}\\ \end{array}\right)\,,\penalty 10000\ \left(\begin{array}[]{c}G_{V}^{\pm}\\ H^{\pm}_{2}\\ \end{array}\right)=\left(\begin{array}[]{cc}c_{\theta_{1}}&-s_{\theta_{1}}\\ s_{\theta_{1}}&c_{\theta_{1}}\\ \end{array}\right)\left(\begin{array}[]{c}\chi_{2}^{\pm}\\ \rho_{3}^{\pm}\\ \end{array}\right)\,$, (59)
where the mixing angles are defined by $t_{\alpha}=\frac{v_{\eta}}{v_{\rho}}$, $t_{\theta_{1}}=\frac{v_{\rho}}{v_{\chi}}$, and $G_{W}^{\pm}$, $G_{V}^{\pm}$ are the Goldstone bosons associated with the longitudinal component of the $W^{\pm}$, $V^{\pm}$, respectively.
In the basis $(I_{\phi},I_{\chi}^{3},I_{\rho}^{2},I_{\eta}^{1})$, the squared mass mixing matrix for the electrically neutral CP odd scalars has form alp331 as below:
${M_{I}^{2}=-\frac{\mathcal{A}}{2}\left(\begin{array}[]{cccc}\frac{1}{v_{\phi}^{2}}&\frac{1}{v_{\phi}v_{\chi}}&\frac{1}{v_{\phi}v_{\rho}}&\frac{1}{v_{\phi}v_{\eta}}\\ &\frac{1}{v_{\chi}^{2}}&\frac{1}{v_{\chi}v_{\rho}}&\frac{1}{v_{\chi}v_{\eta}}\\ &&\frac{1}{v_{\rho}^{2}}&\frac{1}{v_{\eta}v_{\rho}}\\ &&&\frac{1}{v_{\eta}^{2}}\\ \end{array}\right)\,.}$ (64)
As seen from Eq.(64), there are non trivial mixings among the CP-odd scalars $(I_{\phi},I_{\chi}^{3},I_{\rho}^{2},I_{\eta}^{1})$ in the interaction basis.
The CP-odd squared mass mixing matrix $M_{I}^{2}$ in Eq.(64) can be exactly diagonalized alp331 . The relationships between CP-odd scalar fields in the physical states and fields in the interactive states are represented as below:
$a$ $=$ $I_{\phi}c_{\theta_{\phi}}-I_{\chi}^{3}s_{\theta_{3}}s_{\theta_{\phi}}-I_{\rho}^{2}s_{\alpha}c_{\theta_{3}}s_{\theta_{\phi}}-I_{\eta}^{1}c_{\alpha}c_{\theta_{3}}s_{\theta_{\phi}}\,$, (65)
$G_{Z^{\prime}}$ $=$ $I_{\chi}^{3}c_{\theta_{3}}-I_{\rho}^{2}s_{\alpha}s_{\theta_{3}}-I_{\eta}^{1}c_{\alpha}s_{\theta_{3}}\,$, (66)
$G_{Z}$ $=$ $I_{\rho}^{2}c_{\alpha}-I_{\eta}^{1}s_{\alpha}\,$, (67)
$A_{5}$ $=$ $I_{\phi}s_{\theta_{\phi}}+I_{\chi}^{3}s_{\theta_{3}}c_{\theta_{\phi}}+I_{\rho}^{2}s_{\alpha}c_{\theta_{3}}c_{\theta_{\phi}}+I_{\eta}^{1}c_{\alpha}c_{\theta_{3}}c_{\theta_{\phi}}\,$, (68)
with $\alpha$ is given above, two remain angles in this sector are:
$t_{\theta_{3}}=\frac{v_{\eta}}{v_{\chi}}|c_{\alpha}|\,,\hskip 8.5359ptt_{\theta_{\phi}}=\frac{1}{v_{\phi}\sqrt{\frac{1}{v_{\chi}^{2}}+\frac{1}{v_{\rho}^{2}}+\frac{1}{v_{\eta}^{2}}}}\,.$ (69)
Because of the VEVs’ hierachy $v_{\rho},v_{\eta}\ll v_{\chi}\ll v_{\phi}$ as initial assumption, the derived mixing matrix in Eq.(68) has just three angles $\alpha,\theta_{3},\theta_{\phi}$ given and one parameter $\left(\frac{1}{v^{2}_{\phi}}+\frac{1}{v^{2}_{\chi}}+\frac{1}{v^{2}_{\rho}}+\frac{1}{v^{2}_{\eta}}\right)$ which is contained in the expression of $A_{5}$ mass in Eq.(70). Furthermore, the mass of new massive field CP-odd scalar field $A_{5}$} is given by:
$m^{2}_{A_{5}}$ $=$ $-\frac{\mathcal{A}}{2}\left(\frac{1}{v^{2}_{\phi}}+\frac{1}{v^{2}_{\chi}}+\frac{1}{v^{2}_{\rho}}+\frac{1}{v^{2}_{\eta}}\right)\,\approx-\frac{\lambda_{\phi}v_{\phi}v_{\chi}}{s_{2\alpha}}\,.$ (70)
The Eq. (70) shows that the values of $\lambda_{\phi}$ should be negative. It is emphasized that both of the squared mass mixing matrix in Eq. (64) and mass of the $A_{5}$ field are arised from the last term in Eq.(55) which just appears because of the specific discrete symmetry $Z_{11}\otimes Z_{2}$ imposing to this model from the beginning a2 .
Here, the axionlike particle $a$ is massless and given by the combination of four CP odd neutral scalar fields $I_{\phi}$, $I_{\chi}^{3}$, $I_{\rho}^{2}$ and $I_{\eta}^{1}$ as in Eq.(65) which is diffirent from the $a$ field given in Refs.a2 ; jpf . Due to $v_{\chi}\ll v_{\phi}$, the values of $t_{\theta_{\phi}}$ as well as $s_{\theta_{\phi}}$ go to zero, then $c_{\theta_{\phi}}\backsimeq 1$ which approximately helps $a\backsimeq I_{\phi}$.
There are also four fields in the CP-even scalar sector with non-zero VEVs: $R_{\phi},R_{\chi}^{3},R_{\rho}^{2}$ and $R_{\eta}^{1}$. In the basis $(R^{1}_{\eta},R_{\rho}^{2},R_{\chi}^{3},R_{\phi})$, the squared mass mixing matrix of CP-even is defined as alp331 :
$M_{R}^{2}={2}\left(\begin{array}[]{cccc}\lambda_{2}v_{\eta}^{2}-\frac{\mathcal{A}}{4v_{\eta}^{2}}&\frac{\lambda_{6}v_{\eta}v_{\rho}}{4}+\frac{\lambda_{\phi}v_{\chi}v_{\phi}}{4}&\frac{\lambda_{4}v_{\eta}v_{\chi}}{4}+\frac{\lambda_{\phi}v_{\rho}v_{\phi}}{4}&\frac{\lambda_{\phi}v_{\rho}v_{\chi}}{4}+\frac{\lambda_{13}v_{\eta}v_{\phi}}{2}\\ &\lambda_{3}v_{\rho}^{2}-\frac{\mathcal{A}}{4v_{\rho}^{2}}&\frac{\lambda_{\phi}v_{\eta}v_{\phi}}{4}+\frac{\lambda_{5}v_{\rho}v_{\chi}}{2}&\frac{\lambda\phi v_{\eta}v_{\chi}}{4}+\frac{\lambda_{12}v_{\rho}v_{\phi}}{2}\\ &&\lambda_{1}v_{\chi}^{2}-\frac{\mathcal{A}}{4v_{\chi}^{2}}&\frac{\lambda\phi v_{\eta}v_{\rho}}{4}+\frac{\lambda_{11}v_{\chi}v_{\phi}}{2}\\ &&&\lambda_{10}v_{\phi}^{2}-\frac{\mathcal{A}}{4v_{\phi}^{2}}\\ \end{array}\right)\,.$ (75)
With the tiny mixing angles in the CP-even scalar sector are determined as below:
$t_{2\alpha_{2}}=\frac{4c_{\alpha_{3}}v_{\eta}v_{\rho}(\mathcal{A}+\lambda_{6}v_{\eta}^{2}v_{\rho}^{2})}{\mathcal{A}c^{2}_{\alpha_{3}}v_{\eta}^{2}-\mathcal{A}v_{\rho}^{2}+4v_{\eta}^{2}v_{\rho}^{2}(\lambda_{2}v_{\eta}^{2}-\lambda_{3}c^{2}_{\alpha_{3}}v_{\rho}^{2})}\,$, (76)
$t_{2\alpha_{3}}=\frac{4v_{\chi}\left(\mathcal{A}+2\lambda_{5}v_{\rho}^{2}v_{\chi}^{2}\right)}{c_{\alpha_{\phi}}\left(\mathcal{A}-4\lambda_{1}v_{\chi}^{4}\right)^{2}}\,$, (77)
$t_{2\alpha_{\phi}}=\frac{\lambda_{11}v_{\chi}}{\lambda_{10}v_{\phi}}\,$, (78)
the physical fields of CP-even sector are:
$h_{1}$ $=$ $R_{\eta}^{1}c_{\alpha_{2}}+R_{\rho}s_{\alpha_{2}}c_{\alpha_{3}}+R_{\chi}^{3}s_{\alpha_{2}}s_{\alpha_{3}}c_{\alpha_{\phi}}-R_{\phi}s_{\alpha_{2}}s_{\alpha_{3}}s_{\alpha_{\phi}}\,$, (79)
$h_{2}$ $=$ $-R_{\eta}^{1}s_{\alpha_{2}}+R_{\rho}c_{\alpha_{2}}c_{\alpha_{3}}+R_{\chi}^{3}c_{\alpha_{2}}s_{\alpha_{3}}c_{\alpha_{\phi}}-R_{\phi}c_{\alpha_{2}}s_{\alpha_{3}}s_{\alpha_{\phi}}\,$, (80)
$H_{\chi}$ $=$ $-R_{\rho}s_{\alpha_{3}}+R_{\chi}^{3}c_{\alpha_{3}}c_{\alpha_{\phi}}-R_{\phi}c_{\alpha_{3}}s_{\alpha_{\phi}}\,$, (81)
$\Phi$ $=$ $R_{\chi}^{3}s_{\alpha_{\phi}}+R_{\phi}c_{\alpha_{\phi}}\,.$ (82)
In the limit $v_{\phi}\gg v_{\chi}\gg v_{\rho},v_{\eta}$, these physical fields have approximate states and respective masses as below:
$h_{1}$ $\approx$ $R^{1}_{\eta}c_{\alpha_{2}}+R_{\rho}s_{\alpha_{2}}\,,\mbox{\penalty 10000\ with\penalty 10000\ }m_{h_{1},h_{2}}^{2}\approx\lambda_{3}v^{2}+\frac{m_{A_{5}}^{2}}{2}\mp\sqrt{m_{A_{5}}^{4}+\lambda_{3}^{2}\left(v^{4}-3v_{\eta}^{2}v_{\rho}^{2}\right)-\frac{\lambda_{3}m_{A_{5}}^{2}\left(v^{4}-2v_{\eta}^{2}v_{\rho}^{2}\right)}{v^{2}}}\,$,
$h_{2}$ $\approx$ $-R^{1}_{\eta}s_{\alpha_{2}}+R_{\rho}c_{\alpha_{2}}\,$,
$H_{\chi}$ $\approx$ $R^{3}_{\chi}c_{\alpha_{\phi}}\,,\mbox{\penalty 10000\ with\penalty 10000\ }m_{H_{\chi}}^{2}{\approx 2\lambda_{1}v_{\chi}^{2}+\frac{\lambda_{5}^{2}}{2\lambda_{1}}v_{\rho}^{2}}\,$,
$\Phi$ $\approx$ $R_{\phi}c_{\alpha_{\phi}}\,,\mbox{\penalty 10000\ with\penalty 10000\ }m_{\Phi}=\sqrt{2\lambda_{10}v_{\phi}}.\,$ (83)
In comparision with the $4\times 4$ mass mixing matrix of CP-odd sector containing only four parameters with three massless solution, the matrix in Eq.(75) having ten parameters which is unable to be exactly diagonalized. To solve this problem we have used the Hartree-Fock method with conditions such as $v_{\phi}\gg v_{\chi}\gg v_{\rho},v_{\eta}$, $\lambda_{\phi}\ll 1$ and $s_{\alpha_{3}}\approx 0$.
With these VEV hierarchy, the derived matrix contains three angles $\alpha_{2},\alpha_{3}$ and $\alpha_{\phi}$ and three parameters associated with masses of new fields $\Phi,\penalty 10000\ H_{\chi}$ and $h_{2}$. The $\Phi$ field with heavy mass ranging about $10^{11}$ GeV might be used to explain the inflation of the Early Universe alp331 ; Long:2024rdq . The $H_{\chi}$ field is the heavy scalar boson which mass should be at TeV scale. Unfortunately, $\Phi,\penalty 10000\ H_{\chi}$ fields are almost devoid of lepton flavor violating interactions, based on Eqs.(25, 76, 83). Only two lighter Higgs bosons, named as $h_{1},\penalty 10000\ h_{2}$, are involved in lepton flavor violating processes, which we will consider below. Both of these particles have masses at electroweak scale, and the lighter particle, $h_{1}$, is identified with the SM-like Higgs boson.
III Lepton flavor violating decay of charged lepton and CP-even Higgs bosons
In this section, we will be interested in the flavor lepton number violating decays $(l_{a}\rightarrow l_{b}\gamma)$ and $(H\rightarrow l_{a}l_{b}),\,H\equiv h_{1},\,h_{2}$. These channels are closely related, because they share a common source of lepton-flavor violation. Based on known experimental constraints, we will point out suitable regions of the model parameter space where interesting phenomenology and new physics may be found.
III.1 Couplings and Analytical form related to LFV processes
All lepton flavor violating interactions can be given based on the Lagrangian mentioned in Eq.(25), we rewrite them in a convenient form as follows:
$-\mathcal{L}^{Y}_{Lepton}$ $=$ $g_{ab}\overline{\psi}_{aL}l_{bR}\rho+\left(y_{\nu}^{D}\right)_{ab}\overline{\psi}_{aL}\eta N_{bR}+\left(y_{N}\right)_{ab}\phi\overline{N}_{aR}^{C}N_{bR}+\mbox{H.c.},$ (84)
$=$ $g_{ab}\left(\overline{\nu}_{aL}\rho^{+}_{1}+\overline{l}_{aL}\rho^{0}_{2}+\overline{(\nu_{R}^{c})}_{aL}\rho^{+}_{3}\right)l_{bR}+\left(y_{N}\right)_{ab}\phi\overline{N}_{aR}^{C}N_{bR}$
$+$ $\left(y_{\nu}^{D}\right)_{ab}\left(\overline{\nu}_{aL}\eta_{1}^{0}+\overline{l}_{aL}\eta_{2}^{-}+\overline{\nu_{R}^{c}}_{aL}\eta_{3}^{0}\right)N_{bR}+\mbox{H.c.}$
The mass mixing matrices of the leptons are given:
$-\mathcal{L}^{mass}_{Lepton}=\frac{g_{ab}v_{\rho}}{\sqrt{2}}\overline{l}_{aL}l_{bR}+\frac{\left(y_{N}\right)_{ab}v_{\phi}}{\sqrt{2}}\overline{N}_{aR}^{C}N_{bR}+\frac{\left(y_{\nu}^{D}\right)_{ab}v_{\eta}}{\sqrt{2}}\overline{\nu}_{aL}N_{bR}+\mbox{H.c.},$ (85)
with the corresponding assignment,
$m_{l_{a}}=\frac{g_{ab}v_{\rho}}{\sqrt{2}}\delta_{ab},\hskip 8.5359ptM^{D}_{\nu}=y^{D}_{\nu}\frac{v_{\eta}}{\sqrt{2}},\hskip 8.5359ptM_{N}=\sqrt{2}\,y_{N}v_{\phi},$ (86)
and
$g_{ab}=\frac{\sqrt{2}m_{l_{a}}}{v_{\rho}}\delta_{ab},\hskip 8.5359pty^{D}_{\nu}=\frac{\sqrt{2}M^{D}_{\nu}}{v_{\eta}},\hskip 8.5359pty_{N}=\frac{M_{N}}{\sqrt{2}v_{\phi}}.$ (87)
Substituting into Eq.(84), we get the corresponding results for the first and second terms on the right side as:
$g_{ab}\overline{\psi}_{aL}l_{bR}\rho$ $=$ $g_{ab}\left(\overline{\nu}_{aL}\rho^{+}_{1}+\overline{l}_{aL}\rho^{0}_{2}+\overline{(\nu_{R}^{c})}_{aL}\rho^{+}_{3}\right)l_{bR}$
$=$ $\frac{\sqrt{2}m_{l_{a}}}{v_{\rho}}\delta_{ab}\left(\overline{\nu}_{aL}\rho^{+}_{1}+\overline{l}_{aL}\rho^{0}_{2}+\overline{(\nu_{R}^{c})}_{aL}\rho^{+}_{3}\right)l_{bR}$
$\supset$ $\frac{\sqrt{2}m_{l_{a}}}{v_{\rho}}\delta_{ab}\left(\overline{\nu}_{aL}l_{bR}s_{\alpha}H^{+}_{1}+\overline{(\nu_{R}^{c})}_{aL}l_{bR}c_{\theta_{1}}H^{+}_{2}-\overline{l}_{aL}l_{bR}\left(h_{2}c_{\alpha_{2}}+h_{1}s_{\alpha_{2}}\right)\right),$
$\left(y_{\nu}^{D}\right)_{ab}\overline{\psi}_{aL}\eta N_{bR}$ $=$ $\frac{\sqrt{2}\left(M^{D}_{\nu}\right)_{ab}}{v_{\eta}}\left(\overline{\nu}_{aL}\eta_{1}^{0}+\overline{l}_{aL}\eta_{2}^{-}+\overline{\nu_{R}^{c}}_{aL}\eta_{3}^{0}\right)N_{bR}$ (88)
$\supset$ $\frac{\sqrt{2}\left(M^{D}_{\nu}\right)_{ab}}{v_{\eta}}\left(-\overline{\nu}_{aL}\left(h_{1}c_{\alpha_{2}}-h_{2}s_{\alpha_{2}}\right)+\overline{l}_{aL}H_{1}^{-}c_{\alpha}\right)N_{bR}.$
In agreement with current experimental observations, charged leptons are assumed to be non-oscillating and active neutrinos generated masses by a type-I seesaw mechanism. We define transformations between the flavor basis $\{e^{\prime}_{aL,R}\}$, $\left\{n^{\prime}_{i}\right\}=\{\nu^{\prime}_{aL},\penalty 10000\ (N^{\prime}_{aR})^{C}\})$ and the mass basis $\{e_{aL,R}\}$, $\left\{n_{i}\right\}=\{\nu_{aL},\penalty 10000\ (N_{aR})^{C}\}$:
$l^{\prime}_{aL}=l_{aL},\penalty 10000\ \penalty 10000\ l^{\prime}_{aR}=l_{aR},\;n^{\prime}_{aL}=U^{\nu}_{ab}n_{bL},\;n^{{}^{\prime}C}_{aL}=U^{\nu*}_{ab}n_{bL}^{C},\text{\,\,where }U^{\nu}\equiv\left(\begin{array}[]{cc}\mathcal{O}&M^{D}_{\nu}\\ (M^{D}_{\nu})^{T}&M_{N}\\ \end{array}\right),$ (91)
with $M^{D}_{\nu}$ and $M_{N}$ have form as Eq.(86).
From the above expansions, we show the lepton-flavor-violating couplings of this model in table. 2.
Vertex Coupling Vertex Coupling
$\bar{\nu}_{a}l_{b}H_{1}^{+},\,\bar{l}_{b}\nu_{a}H_{1}^{-}$ $i\sqrt{2}U^{\nu*}_{ba}\dfrac{m_{l_{b}}}{v_{\rho}}s_{\alpha}P_{R},\,i\sqrt{2}U^{\nu}_{ba}\dfrac{m_{l_{b}}}{v_{\rho}}s_{\alpha}P_{L}$ $\bar{\nu}_{a}l_{b}H_{2}^{+},\penalty 10000\ \nu_{a}\bar{l}_{b}H_{2}^{-}$ $i\sqrt{2}U^{\nu*}_{ba}\dfrac{m_{l_{b}}}{v_{\rho}}c_{\theta_{1}}P_{R}$, $i\sqrt{2}U^{\nu}_{ba}\dfrac{m_{l_{b}}}{v_{\rho}}c_{\theta_{1}}P_{L}$
$\bar{N}_{a}l_{b}H_{1}^{+},\penalty 10000\ N_{a}\bar{l}_{b}H_{1}^{-}$ $i\sqrt{2}\dfrac{(M_{\nu}^{D})^{*}_{ab}}{v_{\eta}}c_{\alpha}P_{L},\penalty 10000\ i\sqrt{2}\dfrac{(M_{\nu}^{D})_{ab}}{v_{\eta}}c_{\alpha}P_{R}$ $\bar{\nu}_{a}N_{b}h_{2},\penalty 10000\ \bar{\nu}_{a}N_{b}h_{1}$ $i\sqrt{2}U^{\nu*}_{bc}\dfrac{(M_{\nu}^{D})_{ab}}{v_{\eta}}s_{\alpha_{2}}P_{L},\penalty 10000\ i\sqrt{2}U^{\nu*}_{bc}\dfrac{(M_{\nu}^{D})_{ab}}{v_{\eta}}c_{\alpha_{2}}P_{L}$
$\bar{l}_{a}l_{a}h_{1},\penalty 10000\ \bar{l}_{a}l_{a}h_{2}$ $\frac{i\sqrt{2}m_{l_{a}}}{v_{\rho}}s_{\alpha_{2}},\penalty 10000\ \frac{i\sqrt{2}m_{l_{a}}}{v_{\rho}}c_{\alpha_{2}}$ $\bar{N}_{a}N_{b}\phi,\penalty 10000\ H_{\chi}H_{2}^{+}V^{\mu-}$ $\frac{(M_{N})_{ab}}{\sqrt{2}v_{\phi}c_{\alpha_{\phi}}},\penalty 10000\ \frac{igs_{\theta_{1}}}{2c_{\alpha_{\phi}}}\left(p_{0}-p_{p}\right)$
$\bar{\nu}_{a}l_{b}W_{\mu}^{+},\penalty 10000\ \bar{\nu}_{a}l_{b}V_{\mu}^{+}$ $\frac{ig}{\sqrt{2}}U^{\nu*}_{ba}\gamma^{\mu}P_{L}$ $\bar{l}_{b}\nu_{a}W_{\mu}^{-},\penalty 10000\ \bar{l}_{b}\nu_{a}V_{\mu}^{-}$ $\frac{ig}{\sqrt{2}}U^{\nu}_{ab}\gamma^{\mu}P_{L}$
$W^{\mu+}W_{\mu}^{-}h_{1},\penalty 10000\ W^{\mu+}W_{\mu}^{-}h_{2}$ $\frac{igm_{W}}{2}\left(c_{\alpha}c_{\alpha_{2}}-s_{\alpha}s_{\alpha_{2}}\right),\penalty 10000\ \frac{igm_{W}}{2}\left(c_{\alpha}s_{\alpha_{2}}-c_{\alpha_{2}}s_{\alpha}\right)$ $Y^{\mu+}Y_{\mu}^{-}h_{1},Y^{\mu+}Y_{\mu}^{-}h_{2}$ $\frac{igm_{W}}{2}c_{\alpha}c_{\alpha_{2}},\penalty 10000\ \frac{igm_{W}}{2}c_{\alpha}s_{\alpha_{2}}$
$h_{1}H_{1}^{+}W^{\mu-},\penalty 10000\ h_{2}H_{1}^{+}W^{\mu-}$ $\frac{ig}{2}(p_{0}-p_{p})\left(c_{\alpha_{2}}s_{\alpha}+c_{\alpha}s_{\alpha_{2}}\right),\penalty 10000\ -\frac{ig}{2}(p_{0}-p_{p})\left(c_{\alpha}c_{\alpha_{2}}+s_{\alpha}s_{\alpha_{2}}\right)$ $h_{1}H_{2}^{+}Y^{\mu-},\penalty 10000\ h_{2}H_{2}^{+}Y^{\mu-}$ $\frac{ig}{2}(p_{0}-p_{p})c_{\alpha_{2}}c_{\theta_{1}},\penalty 10000\ -\frac{ig}{2}(p_{0}-p_{p})c_{\theta_{1}}s_{\alpha_{2}}$
$h_{1}H_{1}^{+}H_{1}^{-},\penalty 10000\ h_{1}H_{2}^{+}H_{2}^{-}$ $i\lambda_{h_{1}H_{1}H_{1}},\penalty 10000\ -i\lambda_{h_{1}H_{2}H_{2}}$ $h_{2}H_{1}^{+}H_{1}^{-},\penalty 10000\ h_{2}H_{2}^{+}H_{2}^{-}$ $i\lambda_{h_{2}H_{1}H_{1}},\penalty 10000\ -i\lambda_{h_{2}H_{2}H_{2}}$
Table 2: Couplings relating with LFV process in 331ALP. All the couplings were only considered in the unitary gauge.
where:
$\lambda_{h_{2}H^{+}_{1}H^{-}_{1}}$ $=-v\left(-c_{\alpha_{2}}\left(\lambda_{6}+\lambda_{9}\right)c_{\alpha}^{3}+s_{\alpha}s_{\alpha_{2}}\left(2\lambda_{2}+\lambda_{9}\right)c_{\alpha}^{2}-c_{\alpha_{2}}s_{\alpha}^{2}\left(2\lambda_{2}+\lambda_{9}\right)c_{\alpha}+s_{\alpha}^{3}s_{\alpha_{2}}\left(\lambda_{6}+\lambda_{9}\right)\right),$
$\lambda_{h_{2}H^{+}_{2}H^{-}_{2}}$ $=-s_{\alpha}s_{\alpha_{2}}v\left(\lambda_{6}c_{\theta_{1}}^{2}+s_{\theta_{1}}^{2}\lambda_{5}\right)+c_{\alpha}c_{\alpha_{2}}v\left(2\lambda_{2}c_{\theta_{1}}^{2}+s_{\theta_{1}}^{2}\left(\lambda_{5}+\lambda_{8}\right)\right)+c_{\theta_{1}}s_{\theta_{1}}\left(c_{\alpha_{2}}v_{\chi}\lambda_{8}+s_{\alpha_{2}}v_{\phi}\lambda_{\phi}\right),$
$\lambda_{h_{1}H^{+}_{1}H^{-}_{1}}$ $=v\left(s_{\alpha_{2}}\left(\lambda_{6}+\lambda_{9}\right)c_{\alpha}^{3}+c_{\alpha_{2}}s_{\alpha}\left(2\lambda_{2}+\lambda_{9}\right)c_{\alpha}^{2}+s_{\alpha}^{2}s_{\alpha_{2}}\left(2\lambda_{2}+\lambda_{9}\right)c_{\alpha}+c_{\alpha_{2}}s_{\alpha}^{3}\left(\lambda_{6}+\lambda_{9}\right)\right),$
$\lambda_{h_{1}H^{+}_{2}H^{-}_{2}}$ $=s_{\alpha_{2}}\left(c_{\theta_{1}}s_{\theta_{1}}v_{\chi}\lambda_{8}+c_{\alpha}v\left(2\lambda_{2}c_{\theta_{1}}^{2}+s_{\theta_{1}}^{2}\left(\lambda_{5}+\lambda_{8}\right)\right)\right)+c_{\alpha_{2}}\left(s_{\alpha}v\left(\lambda_{6}c_{\theta_{1}}^{2}+s_{\theta_{1}}^{2}\lambda_{5}\right)-c_{\theta_{1}}s_{\theta_{1}}v_{\phi}\lambda_{\phi}\right).$ (92)
[…]
$\mathcal{D}_{(ab)L}$ $=\mathcal{D}_{(ab)L}\left(m_{\nu_{a}},m_{W}\right)+\mathcal{D}_{(ab)L}\left(m_{\nu_{a}},m_{V}\right)+\mathcal{D}_{(ab)L}\left(m_{\nu_{a}},m_{H^{\pm}_{1,2}}\right)+\mathcal{D}_{(ab)L}\left(m_{N_{a}},m_{H^{\pm}_{1}}\right),$
$\mathcal{D}_{(ab)R}$ $=\mathcal{D}_{(ab)R}\left(m_{\nu_{a}},m_{W}\right)+\mathcal{D}_{(ab)R}\left(m_{\nu_{a}},m_{V}\right)+\mathcal{D}_{(ab)R}\left(m_{\nu_{a}},m_{H^{\pm}_{1,2}}\right)+\mathcal{D}_{(ab)R}\left(m_{N_{a}},m_{H^{\pm}_{1}}\right).$ (93)
The total branching ratios of the cLFV processes can be shown based on Refs. Crivellin:2018qmi ; Thuc:2016qva ; Hung:2021fzb ; Quyet:2023dlo and they have form as:
$\mathrm{Br}^{Total}(l_{a}\rightarrow l_{b}\gamma)\simeq\frac{48\pi^{2}}{G_{F}^{2}}\left(\left|\mathcal{D}_{(ab)R}\right|^{2}+\left|\mathcal{D}_{(ab)L}\right|^{2}\right)\mathrm{Br}(l_{a}\rightarrow l_{b}\overline{\nu_{b}}\nu_{a}),$ (94)
where $G_{F}=g^{2}/(4\sqrt{2}m_{W}^{2})$, and for different charge lepton decays, we use experimental data $\mathrm{Br}(\mu\rightarrow e\overline{\nu_{e}}\nu_{\mu})=100\%,\mathrm{Br}(\tau\rightarrow e\overline{\nu_{e}}\nu_{\tau})=17.82\%,\mathrm{Br}(\tau\rightarrow\mu\overline{\nu_{\mu}}\nu_{\tau})=17.39\%$ as given in Refs.Patrignani:2016xqp ; Tanabashi:2018oca ; Zyla:2020zbs .
The couplings in Tab.2 are given based on the characteristics of the model under consideration. Therefore, we get the following interesting results: $\it{i}$) $\bar{N}_{a}l_{b}H_{2}^{+}=0$ leads to diagram 6) with only $H_{1}^{\pm}$ contributions, $\it{ii}$) $N_{a}$ does not interact with charged gauge bosons so diagram 8) does not give contributions for both $h_{1}$ and $h_{2}$ cases. Therefore, all the one-loop Feynman diagrams of $H\rightarrow l_{a}l_{b},\,\,(H\equiv h_{1},\,h_{2})$ decays are given in Fig.1 and the corresponding amplitudes are also shown in appendix B.
Figure 1: Feynman diagrams at one-loop order of $H\rightarrow l_{a}l_{b}$ decays in the unitary gauge, $H\equiv{h_{1},h_{2}}$
The partial width of $\mathrm{H}\rightarrow l_{a}^{\pm}l_{b}^{\mp}$ is
$\Gamma(\mathrm{H}\rightarrow l_{a}l_{b})\equiv\Gamma(\mathrm{H}\rightarrow l_{a}^{+}l_{b}^{-})+\Gamma(\mathrm{H}\rightarrow l_{a}^{-}l_{b}^{+})=\frac{m_{\mathrm{H}}}{8\pi}\left(|\Delta^{(ab)}_{L}|^{2}+|\Delta^{(ab)}_{R}|^{2}\right).$ (95)
We use the conditions for external momentum as: $p^{2}_{a,b}=m^{2}_{a,b}$,  $(p_{a}+p_{b})^{2}=m^{2}_{\mathrm{H}}$ and $m^{2}_{\mathrm{H}}\gg m^{2}_{a,b}$,  this leads to branching ratio of $\mathrm{H}\rightarrow l_{a}^{\pm}l_{b}^{\mp}$ decays can be given
$\mathrm{Br}(\mathrm{H}\rightarrow l_{a}l_{b})=\Gamma(\mathrm{H}\rightarrow l_{a}l_{b})/\Gamma^{\mathrm{total}}_{\mathrm{H}},$ (96)
We need to know $\Gamma^{\mathrm{total}}_{\mathrm{H}}$, which can be addressed through current experimental data of the SM Higgs boson Hung:2019jue , whose main contributions are:
$\Gamma^{\mathrm{total}}_{\mathrm{H}}$ $=$ $\Gamma(\mathrm{H}\rightarrow b\overline{b})+\Gamma(\mathrm{H}\rightarrow c\overline{c})+\Gamma(\mathrm{H}\rightarrow\tau\overline{\tau})+\Gamma(\mathrm{H}\rightarrow W^{*}W)+\Gamma(\mathrm{H}\rightarrow Z^{*}Z)$ (97)
$+$ $\Gamma(\mathrm{H}\rightarrow gg)+\Gamma(\mathrm{H}\rightarrow\gamma\gamma)+\Gamma(\mathrm{H}\rightarrow Z\gamma),$
and their branching ratios corresponding to the mass of $125.09\penalty 10000\ \penalty 10000\ \mathrm{GeV}$ are given in table 3 Denner:2011mq .
Channel $\tau\overline{\tau}$ $b\overline{b}$ $c\overline{c}$ $W^{+}W^{-}$ $Z^{*}Z$ gggg $\gamma\gamma$ $Z\gamma$
Br($\%$) $6.32$ $57.7$ $2.91$ $21.5$ $2.64$ $8.57$ $0.228$ $0.154$
Table 3: The branching ratios of SM Higgs boson decay with mass of $125.09\penalty 10000\ \penalty 10000\ \mathrm{GeV}$.
We also provide the couplings of CP-even Higgs bosons (H) related to Eq.(97) based on Eq.(25) as in table 4.
Couplings $\overline{\tau}\tau$ $\overline{u_{n}}u_{n}$ $\overline{u_{3}}u_{3}$ $\overline{d_{n}}d_{n}$ $\overline{d_{3}}d_{3}$
$h_{1}$ $i\frac{m_{\tau}}{v_{r}}s_{\alpha_{2}}$ $i\frac{m_{u_{n}}}{v_{r}}s_{\alpha_{2}}$ $i\frac{m_{u_{3}}}{v_{e}}c_{\alpha_{2}}$ $i\frac{m_{d_{n}}}{v_{e}}c_{\alpha_{2}}$ $i\frac{m_{d_{3}}}{v_{r}}s_{\alpha_{2}}$
$h_{2}$ $i\frac{m_{\tau}}{v_{r}}c_{\alpha_{2}}$ $i\frac{m_{u_{n}}}{v_{r}}c_{\alpha_{2}}$ $i\frac{m_{u_{3}}}{v_{e}}s_{\alpha_{2}}$ $i\frac{m_{d_{n}}}{v_{e}}s_{\alpha_{2}}$ $i\frac{m_{d_{3}}}{v_{r}}c_{\alpha_{2}}$
Table 4: The coupling of $H$ with fermions.
Therefore, Eq.(97) is rewritten as Eq.(98), where $g_{HGG}$ is given in table 2 and $g_{Hff}$ as in table 4. While $g^{SM}_{hGG}$ and $g^{SM}_{hff}$ are the corresponding couplings of the SM Higgs boson.
$\Gamma^{\mathrm{total}}_{\mathrm{H}}$ $\simeq$ $\sum_{f}\mathrm{Br}(\mathrm{H}\rightarrow f\overline{f})\left(\frac{g_{Hff}}{g^{SM}_{hff}}\right)^{2}\Gamma^{SM}_{h}+\sum_{G}\mathrm{Br}(\mathrm{H}\rightarrow G^{*}G)\left(\frac{g_{HGG}}{g^{SM}_{hGG}}\right)^{2}\Gamma^{SM}_{h}$ (98)
$+$ $\Gamma(\mathrm{H}\rightarrow gg)+\Gamma(\mathrm{H}\rightarrow\gamma\gamma)+\Gamma(\mathrm{H}\rightarrow Z\gamma),$
where $G\equiv W^{\pm},\penalty 10000\ Z$, $H\equiv h_{1},\penalty 10000\ h_{2}$ and $f\equiv\tau,b,c$.
The last three terms of Eq.(98) are loop-induced decays when considered in SM, but $h_{1,2}$ do not interact with exotic quarks so $\Gamma(\mathrm{H}\rightarrow gg)\simeq\Gamma(\mathrm{H}\rightarrow\overline{u_{3}}u_{3}),\penalty 10000\ u_{3}\equiv t-quark$, the remaining two terms are given as,
$\Gamma(\mathrm{H}\rightarrow\gamma\gamma)=\mathrm{Br}(\mathrm{H}\rightarrow\gamma\gamma)\Gamma^{\mathrm{total}}_{\mathrm{H}},\penalty 10000\ \Gamma(\mathrm{H}\rightarrow Z\gamma)=\mathrm{Br}(\mathrm{H}\rightarrow Z\gamma)\Gamma^{\mathrm{total}}_{\mathrm{H}}.$ (99)
By replacing Eq.(99) with Eq.(98) and combining them with table 3 and table 4, we can determine the total decay width of $h_{1}$ and $h_{2}$. Note that in Eq.(98), $\Gamma^{\mathrm{SM}}_{h}\simeq 4.1\times 10^{-3}\penalty 10000\ \mathrm{GeV}$ corresponding to $m_{h}=125.09\,[\mathrm{GeV}]$ CMS:2012hfp ; CMS:2012zhx ; Barger:2012hv and the upper bound of the current experimental limit for $\mathrm{Br}(\mathrm{h}\rightarrow\mu\tau)$ is $10^{-3}$ as shown in Refs. Patrignani:2016xqp ; Tanabashi:2018oca ; Zyla:2020zbs .
III.2 Numerical results
We use the well-known experimental parameters Zyla:2020zbs ; Patrignani:2016xqp : the charged lepton masses $m_{e}=5\times 10^{-4}\,\mathrm{GeV}$,  $m_{\mu}=0.105\,\mathrm{GeV}$,  $m_{\tau}=1.776\,\mathrm{GeV}$,  the SM-like Higgs mass $m_{h}=125.1\,\mathrm{GeV}$,  the mass of the W boson $m_{W}=80.385\,\mathrm{GeV}$, the gauge coupling of the $SU(2)_{L}$ symmetry $g\simeq 0.651$ and $v=246\,\mathrm{GeV}$. The mixing angles act as both free and dependent parameters chosen below to investigate numerical of LFV process. But, a question is difficult as there are so many parameters. This problem can be solved by choosing $t_{\alpha}$ and $m_{H^{\pm}_{2}}$ as free parameters. Because, we can express $t_{\alpha_{2}}$ as dependent on $t_{\alpha_{3}}$, $t_{\alpha_{\phi}}$ and the $\mathrm{VeVs}$ based on the relationships in Eq.(76).
$t_{\alpha_{2}}$ $=\frac{1}{4c_{\alpha_{3}}v_{\eta}v_{\rho}\left(\lambda_{6}v_{\eta}v_{\rho}+\lambda_{\phi}v_{\chi}v_{\phi}\right)}\left(4\lambda_{3}c_{\alpha_{3}}^{2}v_{\eta}v_{\rho}^{3}-c_{\alpha_{3}}^{2}\lambda_{\phi}v_{\eta}^{2}v_{\chi}v_{\phi}-4\lambda_{3}v_{\eta}^{3}v_{\rho}+\lambda_{\phi}v_{\rho}^{2}v_{\chi}v_{\phi}\right$.
$\left.-\frac{1}{2}\sqrt{64c_{\alpha_{3}}^{2}v_{\eta}^{2}v_{\rho}^{2}\left(\lambda_{6}v_{\eta}v_{\rho}+\lambda_{\phi}v_{\chi}v_{\phi}\right){}^{2}+\left(-8\lambda_{3}c_{\alpha_{3}}^{2}v_{\eta}v_{\rho}^{3}+2c_{\alpha_{3}}^{2}\lambda_{\phi}v_{\eta}^{2}v_{\chi}v_{\phi}+8\lambda_{3}v_{\eta}^{3}v_{\rho}-2\lambda_{\phi}v_{\rho}^{2}v_{\chi}v_{\phi}\right){}^{2}}\right),$
$t_{\alpha_{3}}$ $=\frac{2\lambda_{1}v_{\chi}^{4}c_{\alpha_{\phi}}+m_{A_{5}}^{2}c_{\alpha}s_{\alpha}v_{\eta}v_{\rho}c_{\alpha_{\phi}}+\sqrt{c_{\alpha_{\phi}}^{2}\left(m_{A_{5}}^{2}c_{\alpha}s_{\alpha}v_{\eta}v_{\rho}+2\lambda_{1}v_{\chi}^{4}\right){}^{2}+16v_{\rho}^{2}v_{\phi}^{2}\left(m_{A_{5}}^{2}c_{\alpha}s_{\alpha}v_{\eta}-\lambda_{5}v_{\rho}v_{\chi}^{2}\right){}^{2}}}{4v_{\rho}v_{\phi}\left(\lambda_{5}v_{\rho}v_{\chi}^{2}-m_{A_{5}}^{2}c_{\alpha}s_{\alpha}v_{\eta}\right)},$
$t_{\alpha_{\phi}}$ $=\frac{\sqrt{\lambda_{11}^{2}v_{\chi}^{2}+\lambda_{10}^{2}v_{\phi}^{2}}-\lambda_{10}v_{\phi}}{\lambda_{11}v_{\chi}},$ (100)
where the VEVs expressed through other parameters such as: the mass of the new charged gauge boson ($m_{V}^{\pm}$), the mixing angle $\alpha$ ($t_{\alpha}=\frac{v_{\eta}}{v_{\rho}}$), in detail, $v_{\chi}=\sqrt{\frac{4m_{V}^{2}}{g^{2}}-v_{\rho}^{2}},\,v_{\eta}=\frac{vt_{a}}{\sqrt{t_{a}^{2}+1}},\,v_{\rho}=\frac{v}{\sqrt{t_{a}^{2}+1}}$. The $\sin{\phi_{i}}$ and $\cos{\phi_{i}}$ functions of arbitrary angles $\phi_{i}$ can be related to $\tan{\phi_{i}}$ by the formulas: $s_{\phi_{i}}=\frac{t_{\phi_{i}}}{\sqrt{t_{\phi_{i}}^{2}+1}},\,c_{\phi_{i}}=\frac{1}{\sqrt{t_{\phi_{i}}^{2}+1}}$. Furthermore, the Yukawa and self-coupling of scalars constants are chosen to be smaller than $\sqrt{4\pi}$ and $4\pi$ respectively, to satisfy the limits of perturbation theory (Ref.Hung:2019jue ).
Moreover, the interaction constants are chosen as dependent parameters based on Eqs.(58,76), the rest will be treated as free parameters and must be in the range of values that ensure the limits of perturbation theory ($\lambda_{i}\leq 4\pi$) . These dependent parameters are listed as in Eq. (III.2).
$\lambda_{8}$ $=\frac{2\left(m_{H_{2}}^{2}v_{\chi}-v^{2}m_{A_{5}}^{2}c_{\alpha}s_{\alpha}v_{\eta}\right)}{v^{2}v_{\rho}v_{\chi}^{2}},$
$\lambda_{9}$ $=\frac{2\left(m_{H_{1}}^{2}-v^{2}m_{A_{5}}^{2}c_{\alpha}s_{\alpha}\right)}{v^{2}v_{\eta}v_{\rho}},$
$\lambda_{\phi}$ $=-\frac{2m_{A_{5}}^{2}c_{\alpha}s_{\alpha}}{v_{\chi}v_{\phi}},$
$\lambda_{3}$ $=\frac{m_{A_{5}}^{2}v_{\eta}^{2}v_{\rho}^{2}-v^{4}m_{A_{5}}^{2}+m_{h}^{2}v^{4}+\sqrt{\left(m_{A_{5}}^{2}\left(v_{\eta}^{2}v_{\rho}^{2}-v^{4}\right)+m_{h}^{2}v^{4}\right){}^{2}+3v^{4}v_{\eta}^{2}v_{\rho}^{2}\left(m_{h}^{2}m_{A_{5}}^{2}+\frac{3}{4}m_{A_{5}}^{4}-m_{h}^{4}\right)}}{3v^{2}v_{\eta}^{2}v_{\rho}^{2}}.$ (101)
To numerically investigate the $l_{a}\rightarrow l_{b}\gamma$ and the LFVHDs, we next consider the parameterization of the mixing matrix $U^{\nu}$ of neutrinos, starting from the original relation of the seesaw mechanism. It has read:
$\hat{M}^{\nu}\equiv diag(m_{\nu_{1}},m_{\nu_{2}},m_{\nu_{3}})$ $=U^{\dagger}_{PMNS}M^{D}_{\nu}M_{N}^{-1}(M_{\nu}^{D})^{T}U_{PMNS},$ (102)
$U_{PMNS}\hat{M}^{\nu}U^{\dagger}_{PMNS}$ $=M^{D}_{\nu}M_{N}^{-1}(M_{\nu}^{D})^{T}.$ (103)
By using the neutrino oscillation data as shown in Ref.Zyla:2020zbs ; Patrignani:2016xqp ; Tanabashi:2018oca ; ParticleDataGroup:2024cfk ; JUNO:2021vlw : $s^{2}_{12}=0.307,\;s^{2}_{23}=0.558,\;s^{2}_{13}=0.0219,\\ \Delta m^{2}_{21}=7.53\times 10^{-5}\;\mathrm{eV^{2}},\hskip 8.5359pt\Delta m^{2}_{32}=2.455\times 10^{-3}\;\mathrm{eV^{2}}$, and choosing the Dirac and Majorana phases of the $U_{PMNS}$ in their simplest form (Refs.Quyet:2023dlo ; Hung:2022vqx ; Hue:2021zyw ), we can get the right part of Eq.(103) in the known form. Furthermore, $M_{\nu}^{D}$ must be antisymmetric so it contains only six unknowns. Combined with $M_{N}$ in diagonal form, we get three more unknowns. There are a total of nine unknowns on the right part, so Eq.(103) can be solved. That means we have successfully parametrized $U^{\nu}$.
We investigate the numerical results of $l_{a}\rightarrow l_{b}\gamma$ in the parameter space $10^{-3}\leq t_{\alpha}\leq 10^{3}$ and $300\leq m^{\pm}_{H_{2}}\leq 5\times 10^{3}\,\mathrm{GeV}$. We obtain that $\tau\rightarrow e\gamma$ is always smaller than the experimental limit ($4.4\times 10^{-8}$). The part of the parameter space exceeding the upper limit of $\tau\rightarrow\mu\gamma$ is indicated in green, with pink dashed lines corresponding to values $3.3\times 10^{-8}$, $15\times 10^{-8}$ of $\mathrm{Br}(\tau\rightarrow\mu\gamma)$. Meanwhile, the yellow and blue markers are used to represent the parameter space in which $\mathrm{Br}(\mu\rightarrow e\gamma)$ take values larger than the upper bound. We can find the curves for the values $4.2\times 10^{-13}$ and $10\times 10^{-13}$ of $\mathrm{Br}(\mu\rightarrow e\gamma)$ in there. Therefore, the space region allowed to satisfy the experimental constraints on $l_{a}\rightarrow l_{b}\gamma$ decays is the uncolored part in Fig.2.
Figure 2: Contour plots of $l_{a}\rightarrow l_{b}\gamma$ decays in plane of ($t_{\alpha},m_{H^{\pm}_{2}}$).
Based on the analytical form of the amplitudes given in appendix B, we investigate the values of $\mathrm{Br}(h_{1}\rightarrow\mu\tau)$ and $\mathrm{Br}(h_{2}\rightarrow\mu\tau)$, which are given in Fig.3. We can see that, most of the values obtained by $\mathrm{Br}(h_{1}\rightarrow\mu\tau)$, in left panel, and $\mathrm{Br}(h_{1}\rightarrow\mu\tau)$, in right panel, are within the current experimental limits ($\leq 10^{-3}$).
Refer to caption Refer to caption
Figure 3: Plots of $\mathrm{Br}(h_{1}\rightarrow\mu\tau)$ (left panel) and $\mathrm{Br}(h_{2}\rightarrow\mu\tau)$ (right panel) in plane of ($t_{\alpha},m_{H^{\pm}_{2}}$).
To illustrate, we represent the signals of $\mathrm{Br}(h_{1}\rightarrow\mu\tau)$ and $\mathrm{Br}(h_{2}\rightarrow\mu\tau)$, in the allowed space region in Fig.3. On the left panel of Fig.3, $\mathrm{Br}(h_{1}\rightarrow\mu\tau)$ may be get values of $3.4\times 10^{-6}$ and $6.7\times 10^{-6}$ while $\mathrm{Br}(h_{2}\rightarrow\mu\tau)$ may be get values of $9.1\times 10^{-5}$ and $5.6\times 10^{-5}$ in right panel. These values all satisfy the present experimental limits.
Next, we predict the mass of the new CP-even Higgs boson based on guluon-gluon fusion at the LHC and data analysis of CMS. The hadronic cross section depends on the parton luminosity and partonic cross section as follow: Langenfeld:2012ti ; LHCHiggsCrossSectionWorkingGroup:2013rie ,
$\sigma^{h}\left(pp\rightarrow H\right)=\mathcal{L}_{q\overline{q}}\left(m_{H}^{2}\right)\hat{\sigma}\left(\overline{q}_{i}q_{i}\rightarrow H\right).$ (104)
The parton luminosity relate to parton distribution function as:
$\mathcal{L}_{q\overline{q}}\left(m_{H}^{2}\right)=\sum_{q}\int_{\frac{m_{H}^{2}}{s}}^{1}\frac{dx_{1}}{x_{1}}\int_{x_{1}}^{1}\frac{dx_{2}}{x_{2}}\left[F_{q}\left(\frac{x_{1}}{x_{2}},m_{H}^{2}\right)F_{\overline{q}}\left(x_{2},m_{H}^{2}\right)+F_{q}\left(x_{2},m_{H}^{2}\right)F_{\overline{q}}\left(\frac{x_{1}}{x_{2}},m_{H}^{2}\right)\right]\delta\left(x_{1}x_{2}-\frac{m_{H}^{2}}{s}\right),$ (105)
where $F_{q}$ and $F_{\overline{q}}$ are known as the parton distribution functions of quarks and antiquarks, respectively, and $q$ denotes the usual quarks, $\sqrt{s}$ to be the collider center-of mass energy, $\hat{\sigma}$ is the partonic cross section expanded in terms of the strong interaction constant ($\alpha_{s}$) as follows:
[…]
In dilepton decay, if the decay width is assumed to be very narrow then the cross section of these processes can be given approximately as:
$\sigma\left(pp\rightarrow H\rightarrow\overline{l}_{i}l_{j}\right)\simeq\sigma\left(pp\rightarrow H\right)\times\mathrm{Br}\left(H\rightarrow\overline{l}_{i}l_{j}\right).$ (107)
By analyzing the data with center of mass energy at 13 $\mathrm{TeV}$ and integrated luminosity of $35.9\penalty 10000\ \mathrm{fb}^{-1}$ of detector, CMS has shown that $\sigma\left(pp\rightarrow H\right)\times\mathrm{Br}\left(H\rightarrow\mu\tau\right)$ can take values from $51.9\penalty 10000\ (57.4)\penalty 10000\ \mathrm{fb}$ to $1.6\penalty 10000\ (2.1)\penalty 10000\ \mathrm{fb}$ at $95\%$ confidence level CMS:2019pex . To optimize the processing of experimental data, CMS divided into two scales corresponding to low mass 200-450 $\mathrm{GeV}$ and high mass 450-900 $\mathrm{GeV}$.
Figure 4: Plots of $\sigma\left(pp\rightarrow h_{2}\right)\times\mathrm{Br}\left(h_{2}\rightarrow\mu\tau\right)$ (magenta line) depend on $M_{H}$, the green and cyan lines depict the CMS’s observations for the combined decay modes of $H\rightarrow\mu\tau$ (based on Ref. CMS:2019pex ) for the low and high mass ranges, respectively.
As results, we describe the combined two processes $\sigma\left(pp\rightarrow H\right)\times\mathrm{Br}\left(H\rightarrow\mu\tau_{h}\right)$ and $\sigma\left(pp\rightarrow H\right)\times\mathrm{Br}\left(H\rightarrow\mu\tau_{e}\right)$ at $95\%$ CL as the green line and cyan line in Fig. 4, the dotted black represents the boundary between the two mass scales.
To predict the masses of the new CP-even Higgs bosons, $h_{2}$, the parton distribution functions are chosen to be the non-relativistic form (Breit-Wigner distribution functions), the partonic cross section is approximated at the lowest level, and for simplicity $\hat{\sigma}\left(q_{i}\overline{q}_{i}\rightarrow H\right)\sim\sum_{q}g^{2}_{hq\overline{q}}m_{H}\left(1-\frac{4m^{2}_{q}}{m_{H}^{2}}\right)^{3/2}$. So, we investigate signal of $\sigma\left(pp\rightarrow h^{0}_{2}\right)\times\mathrm{Br}\left(h^{0}_{2}\rightarrow\mu\tau\right)$ as magenta line in Fig. 4, and mass of $h_{2}$ is predicted about $600\penalty 10000\ \mathrm{GeV}$.
IV $Z^{\prime}$ boson
In 331ALP, the heavier neutral gauge boson, $Z^{\prime}$, has the mass and physical state given by Eq.(II.2) in section II. The interaction of $Z^{\prime}$ with fermions originates from the kinetic term of spinor fields with the specific form as follows:
$\mathcal{L}^{Z^{\prime}}_{NC}\supset\overline{f}\gamma^{\mu}\left(g_{V}^{Z^{\prime}}(f)+g_{A}^{Z^{\prime}}(f)\gamma^{5}\right)fZ^{\prime}_{\mu},$ (108)
where $g_{V}^{Z^{\prime}}$ and $g_{A}^{Z^{\prime}}$ are given in table 5.
f $\nu_{e},\nu_{\mu},\nu_{\tau}$ $e,\mu,\tau$ $u,c$ $d,s$ $t$ $b$ $U_{3}$ $D_{n}$
$g^{Z^{\prime}}_{V}$ $g\left(\frac{s_{\beta_{1}}}{4\sqrt{3}}-\frac{2tc_{\beta_{1}}}{3\sqrt{6}}\right)$ $g\left(\frac{s_{\beta_{1}}}{4\sqrt{3}}-\frac{2tc_{\beta_{1}}}{3\sqrt{6}}\right)$ $g\left(\frac{s_{\beta_{1}}}{4\sqrt{3}}+\frac{tc_{\beta_{1}}}{3\sqrt{6}}\right)$ $g\left(\frac{s_{\beta_{1}}}{4\sqrt{3}}-\frac{tc_{\beta_{1}}}{6\sqrt{6}}\right)$ $g\left(\frac{s_{\beta_{1}}}{4\sqrt{3}}+\frac{tc_{\beta_{1}}}{2\sqrt{6}}\right)$ $\frac{gs_{\beta_{1}}}{4\sqrt{3}}$ $g\left(-\frac{s_{\beta_{1}}}{2\sqrt{3}}+\frac{tc_{\beta_{1}}}{3\sqrt{6}}\right)$ $\frac{gs_{\beta_{1}}}{2\sqrt{3}}$
$g^{Z^{\prime}}_{A}$ $g\left(\frac{s_{\beta_{1}}}{4\sqrt{3}}-\frac{tc_{\beta_{1}}}{3\sqrt{6}}\right)$ $g\left(\frac{s_{\beta_{1}}}{4\sqrt{3}}-\frac{tc_{\beta_{1}}}{3\sqrt{6}}\right)$ $g\left(\frac{s_{\beta_{1}}}{4\sqrt{3}}+\frac{tc_{\beta_{1}}}{3\sqrt{6}}\right)$ $g\left(\frac{s_{\beta_{1}}}{4\sqrt{3}}-\frac{tc_{\beta_{1}}}{6\sqrt{6}}\right)$ $g\left(\frac{s_{\beta_{1}}}{4\sqrt{3}}+\frac{tc_{\beta_{1}}}{6\sqrt{6}}\right)$ $g\left(\frac{s_{\beta_{1}}}{4\sqrt{3}}-\frac{tc_{\beta_{1}}}{3\sqrt{6}}\right)$ $g\left(-\frac{s_{\beta_{1}}}{2\sqrt{3}}+\frac{tc_{\beta_{1}}}{3\sqrt{6}}\right)$ $g\left(-\frac{s_{\beta_{1}}}{2\sqrt{3}}-\frac{tc_{\beta_{1}}}{3\sqrt{6}}\right)$
Table 5: The coupling of $Z^{\prime}$ with fermions, with $t=\frac{g_{X}}{g}=\frac{3\sqrt{2}t_{W}}{\sqrt{3-t^{2}_{W}}}$ as shown in Ref.Ninh:2005su .
We also use the kinetic energy of the scalar fields, such as Eq.(27), to derive the couplings non-zero of $Z^{\prime}$ with gauge and Higgs bosons as table 6.
Vertex Coupling
$Z^{\prime}H^{+}_{1}H^{-}_{1}$ $\frac{2g\left(c_{\alpha}^{2}\left(\sqrt{4c_{W}^{2}-1}c_{\beta_{1}}\left(\sqrt{3}-3s_{\beta_{1}}\right)-6s_{W}s_{\beta_{1}}s_{\beta_{2}}\right)-s_{\alpha}^{2}\left(\sqrt{4c_{W}^{2}-1}c_{\beta_{1}}\left(3s_{\beta_{1}}+\sqrt{3}\right)+12s_{W}s_{\beta_{1}}s_{\beta_{2}}\right)\right)}{6\sqrt{4c_{W}^{2}-1}}$
$Z^{\prime}H^{+}_{2}H^{-}_{2}$ $\frac{2g\left(2\left(\sqrt{12c_{W}^{2}-3}c_{\beta_{1}}-6s_{W}s_{\beta_{1}}s_{\beta_{2}}\right)c_{\theta_{1}}^{2}+\left(\sqrt{4c_{W}^{2}-1}c_{\beta_{1}}\left(\sqrt{3}-3s_{\beta_{1}}\right)-6s_{W}s_{\beta_{1}}s_{\beta_{2}}\right)s_{\theta_{1}}^{2}\right)}{6\sqrt{4c_{W}^{2}-1}}$
$Z^{\prime}H^{+}_{1}W^{-}$ $\frac{1}{3}gm_{W}c_{\alpha}s_{\alpha}\left(\sqrt{3}c_{\beta_{1}}+\frac{3s_{W}s_{\beta_{1}}s_{\beta_{2}}}{\sqrt{4c_{W}^{2}-1}}\right)$
$Z^{\prime}H^{+}_{2}V^{-}$ $\frac{g^{2}\left(c_{\alpha}c_{\theta_{1}}\left(\sqrt{4c_{W}^{2}-1}c_{\beta_{1}}\left(3s_{\beta_{1}}+\sqrt{3}\right)-24s_{W}s_{\beta_{1}}s_{\beta_{2}}\right)v+\left(\sqrt{4c_{W}^{2}-1}c_{\beta_{1}}\left(3s_{\beta_{1}}+\sqrt{3}\right)+12s_{W}s_{\beta_{1}}s_{\beta_{2}}\right)s_{\theta_{1}}v_{\chi}\right)}{12\sqrt{4c_{W}^{2}-1}}$
Table 6: The coupling of $Z^{\prime}$ with charged and gauge bosons.
The experimental LHC, with the final product of the dilepton at $\sqrt{s}=13\penalty 10000\ \mathrm{TeV}$, suggests that the process in which $Z^{\prime}$ is mediated has a cross section as the product of the hadronic cross section and the branching ratio CMS:2021ctt .
$\sigma\left(pp\rightarrow Z^{\prime}\rightarrow\overline{l}l\right)\simeq\sum_{q}\int_{\tau}^{1}\frac{dx_{1}}{x_{1}}\int_{\tau}^{1}\frac{dx_{2}}{x_{2}}\xi_{q}\left(\frac{x_{1}}{x_{2}},m_{Z}^{2}\right)\overline{\xi}_{q}\left(x_{2},m_{Z}^{2}\right)\hat{\sigma}\left(\overline{q}qi\rightarrow Z\right)\times\mathrm{Br}\left(Z\rightarrow\overline{l}l\right),$ (109)
where $\xi_{q}$ and $\overline{\xi}_{q}$ are the parton distribution functions of quarks and antiquarks, respectively. The partonic cross section and the branching ratio are given as:
$\hat{\sigma}\left(\overline{q}q\rightarrow Z^{\prime}\right)$ $=$ $\frac{\pi g^{2}}{3c^{2}_{W}}\sum_{i}\left[\left(g^{Z^{\prime}}_{V}(q_{i})\right)^{2}+\left(g^{Z^{\prime}}_{A}(q_{i})\right)^{2}\right],$
$\mathrm{Br}\left(Z^{\prime}\rightarrow\overline{l}l\right)$ $=$ $\frac{\Gamma\left(Z^{\prime}\rightarrow\overline{l}l\right)}{\Gamma^{tot}_{Z^{\prime}}},$ (110)
with the decay channel products being dileptons, the decay width has the form as in Eq.(111 ), while the total decay width has to take into account the gauge and Higgs boson contributions, denoted as $G$ and $H$, as listed in Eq.(112).
$\Gamma\left(Z^{\prime}\rightarrow\overline{l}l\right)\simeq\frac{m_{Z^{\prime}}}{12\pi}\frac{\pi g^{2}}{c^{2}_{W}}\sum_{e,\mu,\tau,\nu,N_{a}}\sqrt{1-\frac{4m^{2}_{l}}{m^{2}_{Z^{\prime}}}}\left[\left(g^{Z^{\prime}}_{V}(l)\right)^{2}\left(1+\frac{2m^{2}_{l}}{m^{2}_{Z^{\prime}}}\right)+\left(g^{Z^{\prime}}_{A}(l)\right)^{2}\left(1-\frac{4m^{2}_{l}}{m^{2}_{Z^{\prime}}}\right)\right],$ (111)
$\Gamma^{tot}_{Z^{\prime}}\simeq\Gamma\left(Z^{\prime}\rightarrow\overline{f}f\right)+\Gamma\left(Z^{\prime}\rightarrow H^{*}H\right)+\Gamma\left(Z^{\prime}\rightarrow GH\right)+\Gamma\left(Z^{\prime}\rightarrow G^{*}G\right).$ (112)
Because $Z$ and $Z^{\prime}$ do not mix as shown in Eq.(II.2), $\Gamma(Z^{\prime}\rightarrow W^{+}W^{-})$ is canceled out according to Ref.Perez:2004jc . Furthermore, $m_{X}>m_{Z^{\prime}}/2$ means $\Gamma(Z^{\prime}\rightarrow X^{*}X,\penalty 10000\ X\equiv V^{\pm},X^{0*},X^{0}$ also do not contribute to the total width of $Z^{\prime}$. The remaining components are in the form as follows Ninh:2005su ; cl ; Perez:2004jc :
$\Gamma\left(Z^{\prime}\rightarrow\overline{f}f\right)\simeq\frac{m_{Z^{\prime}}}{12\pi}\frac{\pi g^{2}}{c^{2}_{W}}\sum_{f}N_{C}(f)\sqrt{1-\frac{4m^{2}_{f}}{m^{2}_{Z^{\prime}}}}\left[\left(g^{Z^{\prime}}_{V}(f)\right)^{2}\left(1+\frac{2m^{2}_{f}}{m^{2}_{Z^{\prime}}}\right)+\left(g^{Z^{\prime}}_{A}(f)\right)^{2}\left(1-\frac{4m^{2}_{f}}{m^{2}_{Z^{\prime}}}\right)\right].$ (113)
$\Gamma\left(Z^{\prime}\rightarrow H^{*}H\right)\simeq\frac{m_{Z^{\prime}}}{48\pi}\sum_{H}g_{Z^{\prime}HH}^{2}\left(1-\frac{4m^{2}_{H}}{m^{2}_{Z^{\prime}}}\right)^{3/2}.$ (114)
$\Gamma\left(Z^{\prime}\rightarrow GH\right)\simeq\sum_{G,H}\frac{g_{Z^{\prime}GH}^{2}}{24\pi}\frac{E^{2}-m_{G}^{2}}{m_{Z^{\prime}}^{2}}\left(2+\frac{E^{2}}{m^{2}_{G}}\right),$ (115)
where $E=\frac{m^{2}_{Z^{\prime}}+m^{2}_{G}-m^{2}_{H}}{2m_{Z^{\prime}}}$ is denoted as energy of the gauge boson at the final state and the factors are taken from tables 5 and 6.
Based on the analytical form in Eqs.(109,IV) and the known data, as given in section 14, we show the dependence of $\sigma\left(pp\rightarrow Z^{\prime}\rightarrow\overline{l}l\right)$ on the mass of $m_{Z}$ as blue line in Fig.5. The red line is used to depict the upper limits of dilepton production cross section when ATLAS gives the result at an integrated luminosity of $139\penalty 10000\ \mathrm{fb^{-1}}$ ATLAS:2019erb , while the black line is based on the upper bound given by CMS for $95\%$CL and $\sqrt{s}=13\penalty 10000\ \mathrm{TeV}$ CMS:2021ctt . Therefore, to satisfy the constraints, we can get $m_{Z^{\prime}}\geqslant 5.1\penalty 10000\ \mathrm{TeV}$.
Figure 5: The dependence of $\sigma\left(pp\rightarrow Z^{\prime}\rightarrow\overline{l}l\right)$ on $M_{Z}$ (blue line). The red line represent the upper limits of cross section by ATLAS with $\frac{\Gamma}{m}=6\%$ ATLAS:2019erb , The black line represent the upper limits of cross section by CMS with $\frac{\Gamma}{m}=3\%$ at $95\%$CL CMS:2021ctt
V Dark matter
From the symmetry breaking scheme in (II.1), we see that $Z_{2}$ is the residual symmetry of the model, so odd-$Z_{2}$ particles can be considered as dark matter candidates. These include: $\eta,\,\rho,\,u_{aR},\,d_{nR},\,l_{aR},\,N_{aR}$ according to (4). But $\eta$ and $\rho$ directly break electroweak scale, and $u_{aR},\,d_{nR},\,l_{aR}$ are components that create ordinary matter, so only $N_{aR}$ remains as the best choice to consider as dark matter candidates in the model.
Among $N_{aR}$, we assume $N_{1R}$ is the lightest. So, $N_{1R}$ is considered dark matter and its pair annihilates in to the SM particles in the early universe. Because $N_{1R}$ only direct interacts with $\Phi$ according to Eq.(25), and $\Phi$ can couple with SM particles as Eq.(55), the contribution to the annihilation of dark matter particles mainly comes from the s-channel exchange diagrams mediated by $\Phi$. The diagram is shown in Figure 6.
Figure 6: Dominant contributions to annihilation of dark matter.
Although $\Phi$ mixes with $R_{\chi}$ and $R_{\phi}$ so it can interact with fermions arcooding to Eq.(82), the mixing angles are very small, as Eq.(76). Therefore, the SM product of the dark matter pair annihilation channel does not include fermions but only gauge bosons. It means $N_{1R}^{C}N_{1R}\rightarrow h_{1}h_{1},\penalty 10000\ ZZ$. Hence, product of the annihilation cross-section of the dark matter pair with relative velocity is approximately Cannoni:2015wba ; CMS:2021ctt ,
$\langle\sigma v_{rel}\rangle_{N_{1R}^{C}N_{1R}\rightarrow all}$ $=2\sigma_{\Lambda}F_{-4}(x_{f})\simeq\frac{3}{x_{f}}\sigma_{\Lambda}$ (116)
$\simeq\frac{3m^{2}_{N_{1R}}}{4\pi x_{f}}\left\{\frac{\left(g_{\Phi N_{1R}^{C}N_{1R}}\right)^{2}}{4m^{2}_{N_{1R}}-m^{2}_{\Phi}}\times\left[\left(g_{\Phi ZZ}\right)^{2}+\left(g_{\Phi h_{1}h_{1}}\right)^{2}\right]\right\},$
where $x_{f}\simeq 20.25$ Bertone:2004pz and $g_{\Phi N_{1R}^{C}N_{1R}}=\left(y_{N}\right)_{11}$, while on the contrary $g_{\Phi ZZ}=0$, $g_{\Phi h_{1}h_{1}}=\frac{v_{\phi}}{2\sqrt{2}}\left(\lambda_{12}s_{\alpha_{2}}^{2}+\lambda_{13}c_{\alpha_{2}}^{2}\right)$ are determined from Eq.(27) and Eq.(55), respectively.
We employed the relic abundance of dark matter as in Ref.Planck:2018vyg follows:
$\Omega_{DM}h^{2}\simeq\frac{0.1\penalty 10000\ \mathrm{pb}}{\langle\sigma v_{rel}\rangle_{N_{1R}^{C}N_{1R}\rightarrow all}}.$ (117)
Figure 7: Contour of the dark matter relic density on plane of ($v_{\Phi},\penalty 10000\ m_{N_{1R}}$).
We represent the dark matter relic density, $0.1\leq\Omega_{DM}h^{2}\leq 0.3$ Bahcall:1999xn , in the plane of ($v_{\Phi},\penalty 10000\ m_{N_{1R}}$) as shown in Figure 7. As a result, the colorless is the parameter space region that satisfies current observations, in which there is a relationship between the mass of dark matter and the breaking scale of axion.
VI Conclusions
In the 3-3-1 model with axionlike particles (ALP), we show the region of the parameter space where both cLFV and LFVHDs are satisfied experimental limits. These are allowed parameter spaces where $\mathrm{Br}(h_{1}\rightarrow\mu\tau)$ has a value about of $10^{-6}$ and $\mathrm{Br}(h_{2}\rightarrow\mu\tau)$ reaches a value approximately of $10^{-5}$. Based on gluon-gluon fusion at the LHC, we investigate the signals of cross-sections in the allowed parameter space, and provide predictions of $m_{h_{2}}\geq 600\penalty 10000\ \mathrm{GeV}$. Based on the search for high-mass dilepton resonances at ATLAS and CMS, we take the numerical result of $\sigma\left(pp\rightarrow Z^{\prime}\rightarrow\overline{l}l\right)$ and show that the mass of $Z^{\prime}$-boson is $m_{Z^{\prime}}\geq 5.1\penalty 10000\ \mathrm{TeV}$. The hypothesis of the existence of residual symmetry $Z_{2}$ after spontaneous symmetry breaking stages allows that only $N_{aR}$ are dark matter candidates among the odd -$Z_{2}$ particles. Investigating the relic density of dark matter within experimentally permissible limits, we also established a relationship between the mass of dark matter and the breaking scale of axion.
Acknowledgments
H.T. Hung would like to thank A.V. Bednyakov for the helpful comments to improving this article.
Appendix A Master integrals.
In this section, we use the Passarino-Veltman (PV) functions as mentioned in Ref. Passarino:1978jh to establish the contributions at one-loop order of the Feynman diagrams in Fig. 1. By introducing the formulas $D_{0}=k^{2}-M_{0}^{2}+i\delta$, $D_{1}=(k-p_{1})^{2}-M_{1}^{2}+i\delta$ and $D_{2}=(k+p_{2})^{2}-M_{2}^{2}+i\delta$, with $\delta$ is infinitesimally a positive real quantity, we give:
$A_{0}(M_{n})$ $\equiv$ $\frac{\left(2\pi\mu\right)^{4-D}}{i\pi^{2}}\int\frac{d^{D}k}{D_{n}},\hskip 8.5359ptB^{(12)}_{0}\equiv\frac{\left(2\pi\mu\right)^{4-D}}{i\pi^{2}}\int\frac{d^{D}k}{D_{1}D_{2}},\hskip 8.5359ptB^{(1)}_{0}\equiv\frac{\left(2\pi\mu\right)^{4-D}}{i\pi^{2}}\int\frac{d^{D}k}{D_{0}D_{1}},$
$B^{(2)}_{0} \displaystyle\equiv$ $\frac{\left(2\pi\mu\right)^{4-D}}{i\pi^{2}}\int\frac{d^{D}k}{D_{0}D_{2}},\hskip 8.5359ptC_{0}\equiv C_{0}(M_{0},M_{1},M_{2})=\frac{1}{i\pi^{2}}\int\frac{d^{4}k}{D_{0}D_{1}D_{2}},$ (118)
where $n=1,2$, $D=4-2\epsilon\leq 4$ is the dimension of the integral, while $\penalty 10000\ M_{0},\penalty 10000\ M_{1},\penalty 10000\ M_{2}$ indicate the masses of virtual particles in the loops. We also suppose $p^{2}_{1}=m^{2}_{1},\penalty 10000\ p^{2}_{2}=m^{2}_{2}$ for external particle. The tensor integrals are
$A^{\mu}(p_{n};M_{n})$ $=$ $-\frac{i\left(2\pi\mu\right)^{4-D}}{\pi^{2}}\int\frac{k^{\mu}\times d^{D}k}{D_{n}}=A_{0}(M_{n})p_{n}^{\mu},$
$B^{\mu}(p_{n};M_{0},M_{n})$ $=$ $-\frac{i\left(2\pi\mu\right)^{4-D}}{i\pi^{2}}\int\frac{k^{\mu}\times d^{D}k}{D_{0}D_{n}}\equiv B^{(n)}_{1}p^{\mu}_{n},$
$B^{\mu}(p_{1},p_{2};M_{1},M_{2})$ $=$ $-\frac{i\left(2\pi\mu\right)^{4-D}}{i\pi^{2}}\int\frac{k^{\mu}\times d^{D}k}{D_{1}D_{2}}\equiv B^{(12)}_{1}p^{\mu}_{1}+B^{(12)}_{2}p^{\mu}_{2},$
$C^{\mu}(M_{0},M_{1},M_{2})$ $=$ $-\frac{i}{\pi^{2}}\int\frac{k^{\mu}\times d^{4}k}{D_{0}D_{1}D_{2}}\equiv C_{1}p_{1}^{\mu}+C_{2}p_{2}^{\mu},$
$C^{\mu\nu}(M_{0},M_{1},M_{2})$ $=$ $-\frac{i}{\pi^{2}}\int\frac{k^{\mu}k^{\nu}\times d^{4}k}{D_{0}D_{1}D_{2}}\equiv C_{00}g^{\mu\nu}+C_{11}p_{1}^{\mu}p_{1}^{\nu}+C_{12}p_{1}^{\mu}p_{2}^{\nu}+C_{21}p_{2}^{\mu}p_{1}^{\nu}+C_{22}p_{2}^{\mu}p_{2}^{\nu},$
where $A_{0}$, $B^{(n)}_{0,1}$, $B^{(12)}_{n}$ and $C_{0,n},C_{mn}$ are PV functions. It is noted that only $A_{0}$, $B^{(n)}_{0,1}$, $B^{(12)}_{n}$ contain the divergence which is denoted as
$\Delta_{\epsilon}\equiv\frac{1}{\epsilon}+\ln 4\pi-\gamma_{E},$ (120)
with $\gamma_{E}$ is the Euler constant.
Using the technique as mentioned in Ref. Hue:2017lak , the divergent parts are shown
$\mathrm{Div}[A_{0}(M_{n})]$ $=$ $M_{n}^{2}\Delta_{\epsilon},\hskip 8.5359pt\mathrm{Div}[B^{(n)}_{0}]=\mathrm{Div}[B^{(12)}_{0}]=\Delta_{\epsilon},$
$\mathrm{Div}[B^{(1)}_{1}$] $=$ $\mathrm{Div}[B^{(12)}_{1}]=\frac{1}{2}\Delta_{\epsilon},\hskip 8.5359pt\mathrm{Div}[B^{(2)}_{1}]=\mathrm{Div}[B^{(12)}_{2}]=-\frac{1}{2}\Delta_{\epsilon}.$ (121)
Therefore, the above PV functions can be written in form:
$A_{0}(M)=M^{2}\Delta_{\epsilon}+a_{0}(M),\,\,B^{(n)}_{0,1}=\mathrm{Div}[B^{(n)}_{0,1}]+b^{(n)}_{0,1},\,\,B^{(12)}_{0,1,2}=\mathrm{Div}[B^{(12)}_{0,1,2}]+b^{(12)}_{0,1,2}$, (122)
where $a_{0}(M),\,\,b^{(n)}_{0,1},\,\,b^{(12)}_{0,1,2}$ are finite parts and have a specific form defined as Ref. Thuc:2016qva ; Hung:2022kmv ; Phan:2016ouz for LFVHDs.
Appendix B Analytic formulas at one-loop order to $l_{a}\rightarrow l_{a}b\gamma$ and $\mathrm{H}\rightarrow l_{i}l_{j}$.
In this section, we will give analytical forms at one-loop order for decays of $l_{a}\rightarrow l_{b}\gamma$ and $\mathrm{H}\rightarrow l_{i}l_{j}$. Based on the couplings mentions in Tab.2, decays of $(l_{a}\rightarrow l_{b}\gamma)$ appear only at the loop level. These contributions include: $\mathcal{D}^{CH}_{L,R}(\nu_{a},W^{\pm}),\mathcal{D}^{CH}_{L,R}(\nu_{a},V^{\pm}),\mathcal{D}^{CH}_{L,R}(\nu_{a},H^{\pm}_{1,2}),\text{and}\,\mathcal{D}^{CH}_{L,R}(N_{a},H_{1}^{\pm})$. They can read Hung:2021fzb ; Hung:2022kmv :
$\mathcal{D}^{CH}_{L,R}(\nu_{a},W^{\pm})$ $=-\frac{eg^{2}}{32\pi^{2}m_{W}^{2}}\sum_{n=1}^{9}U^{\nu*}_{an}U^{\nu}_{bn}F(t_{nW}),$
$\mathcal{D}^{CH}_{L,R}(\nu_{a},Y^{\pm})$ $=-\frac{eg^{2}}{32\pi^{2}m_{Y}^{2}}\sum_{n=1}^{9}U^{\nu*}_{(a+3)n}U^{\nu}_{(b+3)n}F(t_{nY}),$
$\mathcal{D}^{CH}_{L,R}(\nu_{a},H^{\pm}_{1,2})$ $=-\frac{eg^{2}f_{s}}{16\pi^{2}m_{W}^{2}}\sum_{k=1}^{9}\left[\frac{\lambda^{L,s*}_{ak}\lambda^{L,s}_{bk}}{m^{2}_{H^{\pm}_{s}}}\times\frac{1-6t_{ks}+3t^{2}_{ks}+2t^{3}_{ks}-6t^{2}_{ks}\ln(t_{ks})}{12(t_{ks}-1)^{4}}\right$.
$+\left.\frac{m_{n_{k}}\lambda^{L,s*}_{ak}\lambda^{\prime R,s}_{bk}}{m^{2}_{H^{\pm}_{s}}}\times\frac{-1+t_{ks}^{2}-2t_{ks}\ln(t_{ks})}{2(t_{ks}-1)^{3}}\right],$
$\mathcal{D}^{CH}_{L,R}(N_{a},H_{1}^{\pm})$ $=-\frac{eg^{2}f_{s}}{16\pi^{2}m_{W}^{2}}\sum_{k=1}^{9}\left[\frac{\lambda^{L,s*}_{ak}\lambda^{L,s}_{bk}}{m^{2}_{H^{\pm}_{s}}}\times\frac{1-6t_{ks}+3t^{2}_{ks}+2t^{3}_{ks}-6t^{2}_{ks}\ln(t_{ks})}{12(t_{ks}-1)^{4}}\right$.\ (123)
Where, we use symbols
$t_{ks}=\left(\frac{m_{n_{k}}}{m_{H^{\pm}_{s}}}\right)^{2},\penalty 10000\ t_{nW}=\left(\frac{m_{n_{n}}}{m_{W^{\pm}}}\right)^{2},\penalty 10000\ t_{nY}=\left(\frac{m_{n_{n}}}{m_{Y^{\pm}}}\right)^{2}.$ (124)
The one-loop factors of the diagrams in Fig.(1) are given below. We used the same calculation techniques as shown in Thuc:2016qva and the symmetry coefficients of the diagrams are quoted as in Refs.Hue:2010xr ; Dong:2009db . We give $m_{l_{i}}\equiv m_{1}$ and $m_{l_{j}}\equiv m_{2}$.
[…]
- $\left.\left.m_{2}^{2}B^{(2)}_{1}+\left(2m_{V}^{2}+m^{2}_{H}\right)m_{F}^{2}\left(C_{0}-C_{1}\right)\right]\right$. (125)
$\left.-\left(2m_{Y}+\frac{m_{1}^{2}-m_{2}^{2}}{m_{Y}}\right)C_{1}+\left(\frac{m_{1}^{2}-m^{2}_{H}}{m_{Y}}+\frac{m_{2}^{2}m^{2}_{H}}{2m_{Y}^{3}}\right)C_{2}\right\},$
[…]
$+$ $\left.\left.m_{1}^{2}B^{(1)}_{1}+(2m_{V}^{2}+m^{2}_{H})m_{F}^{2}(C_{0}+C_{2})\right]\right$.
$\left.+\left(2m_{Y}-\frac{m_{1}^{2}-m_{2}^{2}}{m_{Y}}\right)C_{2}-\left(\frac{m_{2}^{2}-m^{2}_{H}}{m_{Y}}+\frac{m_{1}^{2}m^{2}_{H}}{m_{Y}^{3}}\right)C_{1}\right\},$
$\mathcal{M}^{(2)}_{L}(a_{1},a_{2},v_{1},v_{2},m_{F},m_{Y},m_{H^{\pm}})$ (128)
$=$ $m_{1}\left\{-\frac{a_{2}}{v_{2}}\frac{m_{F}^{2}}{m_{Y}^{2}}\left(B^{(1)}_{1}-B^{(1)}_{0}\right)+\frac{a_{1}}{v_{1}}m_{2}^{2}\left[2C_{1}-\left(1+\frac{m^{2}_{H^{\pm}}-m^{2}_{H}}{m_{Y}^{2}}\right)C_{2}\right]\right$.
$\left.+\frac{a_{2}}{v_{2}}\frac{m_{F}^{2}}{m_{Y}^{2}}\left[(m^{2}_{H^{\pm}}-m^{2}_{H}+m_{Y}^{2})C_{0}+(m^{2}_{H}-m^{2}_{H^{\pm}}+m_{Y}^{2})C_{1}\right]\right\},$
$\mathcal{M}^{(2)}_{R}(a_{1},a_{2},v_{1},v_{2},m_{F},m_{Y},m_{H^{\pm}})$
[…]
$\left.\left.+2(m^{2}_{H}-m_{2}^{2})C_{1}-\frac{m^{2}_{H^{\pm}}-m^{2}_{H}}{m_{Y}^{2}}\left(m^{2}_{F}C_{0}-m_{1}^{2}C_{1}\right)\right)\right]$
$+\left.+\frac{a_{2}}{v_{2}}\frac{m_{F}^{2}}{m_{Y}^{2}}\left[-2m_{Y}^{2}C_{0}-(m^{2}_{H}-m^{2}_{H^{\pm}}+m_{Y}^{2})C_{2}\right]\right\},$
$\mathcal{M}^{(3)}_{L}(a_{1},a_{2},v_{1},v_{2},m_{F},m_{H^{\pm}},m_{Y})$ (130)
[…]
$\left.\left.+2(m_{1}^{2}-m^{2}_{H})C_{2}+\frac{m^{2}_{H}-m^{2}_{H^{\pm}}}{m_{Y}^{2}}\left(m^{2}_{F}C_{0}+m_{2}^{2}C_{2}\right)\right)\right]$
$\left.+\frac{a_{2}}{v_{2}}m_{F}^{2}\left(-2C_{0}+\frac{m_{Y}^{2}+m^{2}_{H^{\pm}}-m^{2}_{H}}{m_{Y}^{2}}C_{1}\right)\right\},$
$\mathcal{M}^{(3)}_{R}(a_{1},a_{2},v_{1},v_{2},m_{F},m_{H^{\pm}},m_{Y})$
$=$ $m_{2}\left\{\frac{a_{2}}{v_{2}}\frac{m_{F}^{2}}{m_{Y}^{2}}\left(B^{(2)}_{1}+B^{(2)}_{0}\right)+\frac{a_{1}}{v_{1}}m_{1}^{2}\left[-2C_{2}+\left(1+\frac{m^{2}_{H^{\pm}}-m^{2}_{H}}{m_{Y}^{2}}\right)C_{1}\right]\right$.
$\left.+\frac{a_{2}}{v_{2}}\frac{m_{F}^{2}}{m_{Y}^{2}}\left[(m_{Y}^{2}+m^{2}_{H^{\pm}}-m^{2}_{H})C_{0}+(m^{2}_{H^{\pm}}-m^{2}_{H}-m_{Y}^{2})C_{2}\right]\right\}.$
$\mathcal{M}^{(4+5)}_{L}(m_{F},m_{Y})$ $=$ $\frac{-m_{1}m_{2}^{2}}{m_{Y}^{3}(m_{1}^{2}-m_{2}^{2})}\left[\left(2m_{Y}^{2}+m_{F}^{2}\right)\left(B^{(1)}_{1}+B^{(2)}_{1}\right)\right$. (131)
$+$ $\left.m_{1}^{2}B^{(1)}_{1}+m_{2}^{2}B^{(2)}-2m_{F}^{2}\left(B^{(1)}_{0}-B^{(2)}_{0}\right)\right$],
$\mathcal{M}^{(4+5)}_{R}(m_{F},m_{Y})$ $=$ $\frac{m_{1}}{m_{2}}\mathcal{M}^{(4+5)}_{L}(m_{F},m_{Y}),$ (132)
$\mathcal{M}^{(6)}_{L}(a_{1},a_{2},v_{1},v_{2},m_{F},m_{H^{\pm}})$ $=$ $\frac{m_{1}m^{2}_{F}}{v_{2}}\times\left[\frac{a_{1}^{2}}{v_{1}^{2}}m_{2}^{2}(2C_{2}+C_{0})+\frac{a_{2}^{2}}{v_{2}^{2}}m_{F}^{2}(C_{0}-2C_{1})\right$.
$+$ $\left.\frac{a_{1}a_{2}}{v_{1}v_{2}}\left(B^{(12)}_{0}+(m_{F}^{2}+m^{2}_{H^{\pm}}+m_{2}^{2})C_{0}+\frac{}{}2m_{2}^{2}C_{2}-(m_{1}^{2}+m_{2}^{2})C_{1}\right)\right],$
$\mathcal{M}^{(6)}_{R}(a_{1},a_{2},v_{1},v_{2},m_{F},m_{H^{\pm}})$ $=$ $\frac{m_{2}m^{2}_{F}}{v_{2}}\times\left[\dfrac{a_{1}^{2}}{v_{1}^{2}}m_{1}^{2}(C_{0}-2C_{1})+\frac{a_{2}^{2}}{v_{2}^{2}}m_{F}^{2}(C_{0}+2C_{2})\right$.
$+$ $\left.\frac{a_{1}a_{2}}{v_{1}v_{2}}\left(B^{(12)}_{0}+(m_{F}^{2}+m^{2}_{H^{\pm}}+m_{1}^{2})C_{0}-2m_{1}^{2}C_{1}+(m_{1}^{2}+m_{2}^{2})C_{2}\right)\right],$
$\mathcal{M}^{(7)}_{L}(a_{1},a_{2},v_{1},v_{2},m_{F},m_{H^{\pm}})$ $=$ $m_{1}v_{2}\left[\frac{a_{1}a_{2}}{v_{1}v_{2}}m_{F}^{2}C_{0}+\frac{a^{2}_{2}}{v^{2}_{2}}m_{F}^{2}C_{1}-\frac{a^{2}_{1}}{v^{2}_{1}}m_{2}^{2}C_{2}\right],$
$\mathcal{M}^{(7)}_{R}(a_{1},a_{2},v_{1},v_{2},m_{F},m_{H^{\pm}})$ $=$ $m_{2}v_{2}\left[\frac{a_{1}a_{2}}{v_{1}v_{2}}m_{F}^{2}C_{0}+\frac{a^{2}_{1}}{v^{2}_{1}}m_{1}^{2}C_{1}-\frac{a^{2}_{2}}{v^{2}_{2}}m_{F}^{2}C_{2}\right],$
$\mathcal{M}^{(8)}_{L}(m_{Y},m_{F})$ $=$ $\frac{m_{1}m^{2}_{F}}{m_{Y}^{3}}\times\left[\left(B^{(12)}_{0}+B^{(1)}_{1}-(m_{1}^{2}+m_{2}^{2}-2m_{F}^{2})C_{1}\right)-\left(C_{0}-4C_{1}\right)m_{Y}^{2}\right],$
$\mathcal{M}^{(8)}_{R}(m_{Y},m_{F})$ $=$ $\frac{m_{2}m^{2}_{F}}{m_{Y}^{3}}\times\left[\left(B^{(12)}_{0}-B^{(2)}_{1}+(m_{1}^{2}+m_{2}^{2}-2m_{F}^{2})C_{2}\right)-\left(C_{0}+4C_{2}\right)m_{Y}^{2}\right],$
$\mathcal{M}^{(9+10)}_{L}(a_{1},a_{2},v_{1},v_{2},m_{F},m_{H^{\pm}})$ $=$ $\frac{m_{1}}{(m_{1}^{2}-m_{2}^{2})v_{1}}\left[m^{2}_{2}\left(m^{2}_{1}\frac{a_{1}^{2}}{v_{1}^{2}}+m^{2}_{F}\frac{a^{2}_{2}}{v^{2}_{2}}\right)\left(B_{1}^{(1)}+B_{1}^{(2)}\right)\right$. (139)
$\left.\hskip 51.21504pt+m^{2}_{F}\frac{a_{1}a_{2}}{v_{1}v_{2}}\left(2m^{2}_{2}B_{0}^{(1)}-(m^{2}_{1}+m^{2}_{2})B_{0}^{(2)}\right)\right],$
$\mathcal{M}^{(9+10)}_{R}(a_{1},a_{2},v_{1},v_{2},m_{F},m_{H^{\pm}})$ $=$ $\frac{m_{2}}{(m_{1}^{2}-m_{2}^{2})v_{1}}\left[m^{2}_{1}\left(m^{2}_{2}\frac{a_{1}^{2}}{v_{1}^{2}}+m^{2}_{F}\frac{a^{2}_{2}}{v^{2}_{2}}\right)\left(B_{1}^{(1)}+B_{1}^{(2)}\right)\right$. (140)
$\left.\hskip 51.21504pt+m^{2}_{F}\frac{a_{1}a_{2}}{v_{1}v_{2}}\left(-2m^{2}_{1}B_{0}^{(2)}+(m^{2}_{1}+m^{2}_{2})B_{0}^{(1)}\right)\right].$
In the 7th diagram of Fig.(1) we always get that $\mathcal{M}^{(7)}_{L,R}(a_{1},a_{2},v_{1},v_{2},m_{F},m_{H})$ are finite because $C_{k}$ ($k=0,1,2$) do not contain divergent functions.